### v1

# Bounding the conditional von-Neumann entropy for device independent cryptography and randomness extraction

Abstract
This paper introduces a numerical framework for establishing lower bounds on the conditional von-Neumann entropy in device-independent quantum cryptography and randomness extraction scenarios. Leveraging a hierarchy of semidefinite programs derived from the Navascués-Pironio-Acín (NPA) hierarchy, our tool enables efficient computation of entropy bounds based solely on observed statistics, assuming the validity of quantum mechanics. The method’s computational efficiency is ensured by its reliance on projective operators within the non-commutative polynomial optimization problem. The method facilitates provable bounds for extractable randomness in noisy scenarios and aligns with modern entropy accumulation theorems. Consequently, the framework offers an adaptable tool for practical quantum cryptographic protocols, expanding secure communication possibilities in untrusted environments.
I Introduction
In the realm of quantum mechanics, randomness is not merely a consequence of incomplete knowledge but an intrinsic property of nature itself, a principle famously highlighted by Max Born. This inherent randomness has profound implications for quantum information theory, particularly in applications that require the generation of secure random numbers or private cryptographic keys [1, 2]. The quantification of such randomness hinges on the concept of entropy, especially the conditional von-Neumann entropy:
$H(A|Q_{E})_{\rho_{AQ_{E}}}\coloneqq S(\rho_{AQ_{E}})-S(\rho_{Q_{E}}).$ (1)
Here, $\rho_{AQ_{E}}$ represents a classical-quantum state in the post-measurement system $A$ on Alice’s side, with quantum side-information $Q_{E}$, which we assume is fully accessible to an eavesdropper. The conditional von-Neumann entropy measures the maximum randomness extractable in a secure, asymptotic manner from quantum systems under certain constraints. Given its importance, lower bounding the conditional von-Neumann entropy has become a central task in both theoretical studies and practical applications of quantum information and especially security proofs in quantum cryptography [3, 4, 5].
Recent advancements in quantum cryptography, particularly in device-independent quantum key distribution (DIQKD) and randomness extraction [6, 7, 8, 9, 10], harness the intrinsic unpredictability of quantum systems to secure communication against eavesdropping adversaries. DIQKD protocols rely on the principle that correlations observed in measurement outcomes from entangled quantum systems cannot be accounted for by any local hidden variable theory, as shown by violations of Bell inequalities [11, 12, 13, 14]. This fundamental characteristic implies that a conditional distribution
$p(a,b|x,y)$ (2)
cannot be reproduced if $a$ and $b$ were deterministic outputs in hands of an evesdropper for each input pair $(x,y)$, with values governed by a shared random variable, like a common coin, which randomizes $a$ and $b$ jointly [15]. Instead, quantum non-locality, revealed through such correlations, ensures that outcomes cannot be explained by classical models, thus establishing true randomness by laws of quantum theory. The DIQKD approach offers an additional layer of security by requiring no trust in the measurement devices themselves. Given this framework, device-independent protocols make the accurate bounding of entropy under non-asymptotic conditions a central challenge [16, 17, 18]. The need to rigorously quantify the randomness of measurement outcomes, especially in real-world scenarios where noise and finite sample sizes introduce deviations, underscores the importance of non-asymptotic entropy bounds. Such bounds provide critical insights into the security achievable in device-independent settings, where only the observed distribution and the validity of quantum theory are assumed [19, 20, 21, 22, 23]. As research advances, these entropy bounds are becoming essential for verifying and implementing DIQKD protocols in practical cryptographic systems, thus expanding the possibilities of secure communication in untrusted environments [24, 25, 26].
In this work, we contribute to the field of device-independent randomness extraction, addressing how much usable randomness can be extracted from a given experiment. We introduce a new numerical method that provides provable lower bounds for the extractable randomness in (1). Thus, we answer the question of how much randomness can be extracted from a specific type of experiment, assuming only the availability of measurement results from Alice and Bob and the premise that the experiment is governed by quantum theory. Our method complements the existing work of Brown, Fawzi, and Fawzi [3], as well as Tan et al. [4]. The demand for such tools has increased recently due to significant progress in entropy accumulation theorems [19, 20, 21, 22, 23], which provide concrete solutions for randomness extraction in finite-size regimes by reducing it to the asymptotic quantity in (1). Thus, after applying an entropy accumulation theorem, the remaining task is to calculate provable lower bounds for the conditional von-Neumann entropy (1) for specific noise models and input-output data scenarios. Furthermore, by confining the analysis to projective operators, our method achieves a high degree of computational efficiency, making it suitable for real-world applications.
II Bounding the von-Neumann entropy
[…]
$D(\rho\|\sigma)\coloneqq\begin{cases}\operatorname{tr}[\rho\log\rho-\rho\log\sigma]&\ker[\sigma]\subseteq\ker[\rho]\\ +\infty&\text{else}\end{cases}$ (3)
by P. E. Frenkel [27] for two arbitrary positive operators $\sigma,\rho\in\mathcal{P}(\mathcal{H})$ on a Hilbert space $\mathcal{H}$
[…]
For the rest of the work we use the expression $\operatorname{tr}^{-}[A]$ for the trace of the negative part of a self-adjoint operator $A$ on $\mathcal{H}$ with unique Jordan decomposition $A=A_{+}-A_{-}$ and positive operators $A_{+}$ respectively $A_{-}$. Similarly we define $\operatorname{tr}^{+}[A]$. As shown by Jenčová [28], the formula in (4) is even for infinite-dimensional separable Hilbert spaces true and can be rewritten in the following simple form
$D(\rho\|\sigma)=\frac{1}{\ln 2}\big{(}\operatorname{tr}[\rho-\sigma]$ $+\int_{\mu}^{\lambda}\frac{ds}{s}\operatorname{tr}^{+}[\sigma s-\rho]$ (5)
$+\ln\lambda+1-\lambda\big{)}.$
The integral bounds $\mu,\lambda\in\mathbb{R}_{\geq 0}$ are solutions for the operator valued inequality
$\mu\sigma\leq\rho\leq\lambda\sigma.$ (6)
The value of $\lambda$ can be found by solving the equation $\log\lambda=D_{\operatorname{max}}(\rho\|\sigma)$, which defines the max-relative entropy, $D_{\operatorname{max}}$ [29]. For the remainder of the work, we assume a tripartite experiment involving Alice, Bob, and Eve, modeled by the axioms of quantum theory as a tripartite state, a positive operator on a Hilbert space with trace equal to $1$,
$\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{QE}).$ (7)
Importantly, the argument for the device independence specifies nothing concrete about the Hilbert space $\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{QE}$. In the following we follow the example of generating a raw-key in a DIQKD protocol, even though our methods can be applied for a larger class of problems section III. Afterwards, we discuss briefly some simple generalizations.
The fact that we consider the von-Neumann entropy means that we do not have to deal with finite size effects. In section III we argue this assumption for certain QKD scenarios. In a DIQKD protocol we usually have some measurements $\mathcal{X}$ on $\mathcal{S}(\mathcal{H}_{Q_{A}})$ with finite outcome set $A$ and similarly $\mathcal{Y}$ on $\mathcal{S}(\mathcal{H}_{Q_{B}})$ with finite outcome set $B$ respectively. Measurements are generally understood as sets of POVMs $\{M_{a|x}\}_{a\in A}$ for each $x\in\mathcal{X}$ and similarly $\{M_{b|y}\}_{b\in B}$ for each $y\in\mathcal{Y}$. In a more abstract language we understand it as a quantum channel from a commutative algebra into our observable algebra. In the most simplest form of a protocol, the raw-key is now constructed by one specific measurement $\tilde{x}\in\mathcal{X}$. This can be generalized later on [30], but we skip the technical details of it for now. In a next step, we use the fact that the asymptotically extractable randomness out of the fixed measurement $\tilde{x}\in\mathcal{X}$ can be described as [^2] For simplicity we abbreviate $\rho_{AQ_{E}|\tilde{x}}\equiv\rho_{AQ_{E}|X=\tilde{x}}$
$H(A|X=\tilde{x},Q_{E})_{\rho_{AQ_{E}|\tilde{x}}}=-D(\rho_{AQ_{E}|\tilde{x}}\|\mathds{1}_{A}\otimes\rho_{Q_{E}}).$ (8)
This yields that we are particular interested in lower bounding $-D(\rho_{AQ_{E}|\tilde{x}}\|\mathds{1}_{A}\otimes\rho_{Q_{E}})$. Specifically, device independent lower bounds satisfy the following prototypical type of optimization problem
$\inf$ $-D(\rho_{AQ_{E}|\tilde{x}}\|\mathds{1}_{A}\otimes\rho_{Q_{E}})$ (9)
$\sum_{abxyi}c_{abxyi}\operatorname{tr}[M_{a|x}\otimes N_{b|y}\rho_{Q_{A}Q_{B}}]\geq q_{i},\quad 1\leq i\leq m$
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X}$
$\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X}$
$N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
$\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}})$
whereby we do not aim to specify the concrete underlying Hilbert space. Moreover, the second line can be understood as information about the joint statistics of Alice and Bob, which is nothing more than (2). A simple but sufficient and common example would be that the $c_{abxyi}\in\mathbb{R}$ coefficients form a Bell inequality [31, 3, 18]. In the following we need to estimate the program (9) with an actually quantifiable program. For this purpose, we use the integral representation (5) and derive an upper bound for the relative entropy (3) with an operator valued optimization problem.
Theorem II.1 (Upper bounds for relative entropy).
Consider two positive operators $\rho,\sigma\in\mathcal{P}(\mathcal{H})$ such that
$\mu\sigma\leq\rho\leq\lambda\sigma$ (10)
for arbitrary $\mu,\lambda\in\mathbb{R}_{\geq 0}$. Then we can upper bound
$\int_{0}^{\lambda}\frac{ds}{s}\operatorname{tr}^{+}[\sigma s-\rho]\leq\sup_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\sum_{k=0}^{r}\operatorname{tr}[P_{k}(\alpha_{k}\rho+\beta_{k}\sigma)]$ (11)
for sets of real numbers $\alpha_{0},\ldots,\alpha_{r},\beta_{0},\ldots,\beta_{r}\in\mathbb{R}$, which can be easily precomputed.
In particular the estimate (11) yields
$D(\rho\|\sigma)$ $\leq\frac{1}{\ln 2}(\operatorname{tr}[\rho-\sigma]$ (12)
$+\sup_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\sum_{k=0}^{r}\operatorname{tr}[P_{k}(\alpha_{k}\rho+\beta_{k}\sigma)]$
$+\ln\lambda+1-\lambda).$
Proof.
See Appendix A. ∎
The Theorem II.1 provides a sequence of upper bounds that converge to the relative entropy. Specifically, this sequence of optimization problems can be understood as an application of [32, Appendix E]. We emphasize that a finite $\lambda$ is, at least in finite dimensions, equivalent to the statement that the relative entropy is finite. This is a crucial component in these optimization problems, allowing for effective numerical analysis. In the case of conditional von-Neumann entropy, a finite $\lambda$ always exists, as for any bipartite state $\rho_{AB}$ we have the fundamental operator inequality (see, e.g., [33, Appendix A]):
$\rho_{AB}\leq d_{A}\mathds{1}_{A}\otimes\rho_{B}.$ (13)
Now, let us examine the specific optimization problem in (9). It can be lower-bounded using Theorem II.1 as an optimization over operators $P_{0},\ldots,P_{r}$. In particular, we observe that the optimal operators $P_{k}$, $0\leq k\leq r$, are projections, as they correspond to projections onto the positive part in the Jordan decomposition of the Hermitian operator $\alpha_{k}\rho+\beta_{k}\sigma$. Furthermore, by substituting the specific operators $\rho=\mathds{1}_{A}\otimes\rho_{Q_{E}}$ and $\sigma=\rho_{AQ_{E}|\tilde{x}}$, we find that $\lambda=1$ and $\mu$ can be set to 00 without loss of generality (see Appendix A for details). Thus, we can rewrite (9) as
$\inf$ $\inf_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\frac{1}{\ln 2}(\operatorname{tr}[\mathds{1}_{A}\otimes\rho_{Q_{E}}-\rho_{AQ_{E}|\tilde{x}}]$ (14)
$+$ $\sum_{k=0}^{r}\operatorname{tr}[P_{k}(-\alpha_{k}\rho_{AQ_{E}|\tilde{x}}-\beta_{k}\mathds{1}_{A}\otimes\rho_{Q_{E}})]$
$\sum_{abxyi}c_{abxyi}\operatorname{tr}[M_{a|x}\otimes N_{b|y}\rho_{Q_{A}Q_{B}}]\geq q_{i},\quad 1\leq i\leq m$
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X},\quad\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X},\quad N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
$\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}})$
$P_{k}^{2}=P_{k}\in\mathcal{P}(\mathcal{H}_{A}\otimes\mathcal{H}_{Q_{E}}).$
An application of Naimark’s dilation theorem (e.g., [34, Chap. Dilation Theory]) additionally allows us to rewrite the state $\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}})$ without loss of generality as a pure state $\psi_{Q_{A}Q_{B}Q_{E}}\in\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}}$, with all measurements also being projective. By carefully examining the classical-quantum structure of $\rho_{AQ_{E}|\tilde{x}}$, the post-measurement state in our specific key basis $\tilde{x}\in\mathcal{X}$ can then be used to reformulate (14) as a non-commutative polynomial optimization in Theorem II.2.
Theorem II.2 (NPO Formulation).
Calculating lower bounds for the conditional von-Neumann entropy (1) in the device independent setting with knowledge about statistics between Alice and Bobs systems can be reformulated into the following non-commutative polynomial optimization problem
$H(A|$ $X=\tilde{x},Q_{E})\geq$ (15)
$\inf$ $\frac{1}{\ln 2}(|A|-1+$
$\sum_{k=0}^{r}\sum_{a\in A}\langle\psi_{Q_{A}Q_{B}Q_{E}}|-\alpha_{k}M_{a|x}P_{k}^{(a)}$
$-\beta_{k}P_{k}^{(a)}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle)$
$\sum_{abxyi}c_{abxyi}\langle\psi_{Q_{A}Q_{B}Q_{E}}|M_{a|x}N_{b|y}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle\geq q_{i}$
$\quad 1\leq i\leq m$
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X},\quad\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X},\quad N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
$[M_{a|x},N_{b|y}]=[M_{a|x},P_{k}^{(a)}]=[N_{b|y},P_{k}^{(a)}]=0$
[…]
$(P_{k}^{(a)})^{2}=P_{k}^{(a)},\quad(P_{k}^{(a)})^{\star}=P_{k}^{(a)}$
$\quad 1\leq k\leq r,\ a\in A.$
Proof.
See Appendix B
The polynomial optimization problem (15) can be interpreted as an optimization over the algebraic constraints within it. In a finite-dimensional Hilbert space setting, these constraints can be viewed as generating rules of a $\star$-subalgebra of $\mathcal{B}(\mathcal{H})$, where $\mathcal{H}$ is a Hilbert space (in infinite dimensions, caution is required to handle the topological closure of the algebra, though this remains feasible). Assuming that these relations form an algebra $\mathcal{A}$, we employ the Navascues-Pironio-Acin (NPA) hierarchy [35], applied in the algebraic framework as demonstrated in [36], to compute outer approximations to the optimization problem (15), thus lower bounds.
The approach involves relaxing the positivity requirement for linear, continuous functionals on $\mathcal{A}$. Rather than requiring positivity across the positive cone $\mathcal{A}_{+}$, we restrict it to a sum-of-squares cone $\Sigma_{2}^{(n)}\subset\mathcal{A}_{+}$ at level $n$. Optimization over this relaxed SDP can then be viewed as optimizing the dual cone $(\Sigma_{2}^{(n)})^{\star}$, which is representable in SDP form and provides a convergent hierarchy of outer approximations to (15). For further details, we refer to [37, 35, 36, 38] and Appendix D. In our applications, we use the software package from [39].
0.750.750.750.750.80.80.80.80.850.850.850.85000.50.50.50.51111CHSH valueconditional entropylower boundanalytic bound
0.750.750.750.750.80.80.80.80.850.850.850.85000.50.50.50.511111.51.51.51.5CHSH valueconditional entropylower bound $2$
0.80.80.80.80.90.90.90.911110011112222white noiseconditional entropylower bound 2322
Figure 1: The plots illustrate examples of lower bounds for conditional entropies across different scenarios. In the leftmost plot, we present a proof-of-principle example for lower bounds on the conditional entropy $H(A|X=0,Q_{E})_{\rho}$, depending on various values of the CHSH violation in a $2$-scenario (where each of two parties can choose between two measurements with two outcomes). This example is particularly significant because it allows for a verification of the general correctness of the method; as shown in [9], there are analytical values available for this scenario, enabling comparisons with computed results to validate accuracy. The middle plot displays the lower bounds on global randomness, specifically the conditional entropy $H(AB|X=0,Y=0,Q_{E})$, again in a $2$-scenario. Here, the dependency is on the CHSH inequality, and this plot serves to examine the impact of the CHSH violation on the joint conditional entropy of outcomes $A$ and $B$ given measurement settings $X=0$ and $Y=0$. Finally, the rightmost plot demonstrates the quantity $H(AB|X=0,Y=0,Q_{E})$ in a 2322-scenario, where each party has two possible outcomes and Bob has three measurement choices. In this case, we condition on the full distribution of a Werner state depending on dephasing noise.
III Applications and numerical examples
In this section, we present numerical evidence supporting the methods and tools developed in section II. To structure this discussion, we divide section III into two parts, as outlined in [5].
In the first part, we examine a CHSH game scenario with two possible inputs and two possible outputs. Our objective here is to compute $H(A|X=0,Q_{E})$, which is a quantity of particular interest in two main areas: one-sided device-independent randomness extraction (DIRE) and device-independent quantum key distribution (DIQKD). In one-sided DIRE, $H(A|X=0,Q_{E})$ provides a measure of the randomness that can be reliably extracted from Alice’s outcomes, assuming a fixed input $X=0$ and some quantum side information $Q_{E}$. Similarly, in DIQKD, this conditional entropy helps in determining the amount of secure randomness extractable from the quantum correlations between Alice and an adversary.
In the second part, we extend this investigation to the two-sided DIRE task, again within a CHSH game framework. Here, the focus shifts to the quantity $H(AB|X=0,Y=0,Q_{E})$, where the conditional entropy now includes both Alice’s and Bob’s outcomes. This joint entropy becomes essential for evaluating the global randomness accessible when both parties contribute to the measurement outcomes in the context of device-independent randomness extraction.
For clarity, we abbreviate scenarios based on the number of inputs and outputs per party. Specifically, a scenario with $n_{a}$ inputs for Alice, $n_{b}$ inputs for Bob, $o_{a}$ possible outcomes for each input of Alice, and $o_{b}$ possible outcomes for each input of Bob is denoted by $n_{a}n_{b}o_{a}o_{b}$. For instance, the CHSH scenario with two inputs and two outputs per party is a $2$-scenario.
III.1 one-sided randomness
In device-independent quantum information tasks, one-sided (device-independent) randomness is quantified as follows:
$H(A|X=\tilde{x},Q_{E}),$ (16)
where this quantity captures the uncertainty in the outcome $A$ when measurement $X$ takes a fixed value $\tilde{x}$, conditioned on quantum side information $Q_{E}$ held by an eavesdropper. This concept is fundamental in the context of device-independent quantum key distribution (DIQKD) and one-sided randomness extraction [5, 40].
In quantum key distribution (QKD), the asymptotic key rate $r_{\rightarrow}$ can be bounded using the Devetak-Winter formula, which states:
$r_{\rightarrow}\geq H(A|E)_{\rho_{Q_{A}Q_{B}Q_{E}}}-H(A|B)_{\rho_{Q_{A}Q_{B}Q_{E}}},$ (17)
where $H(A|E)_{\rho}$ denotes the conditional entropy of $A$ given the eavesdropper’s knowledge $E$, and $H(A|B)_{\rho}$ denotes the conditional entropy of $A$ given the legitimate party $B$. In DIQKD experiments, estimating $H(A|E)_{\rho}$ relies on observed statistics, such as violations of Bell inequalities or even the full probability distribution gathered from the experiment.
This quantity remains significant even in finite-sample-size regimes due to recent advancements in entropy accumulation theorems (EAT) [19, 20, 21, 22, 23]. These theorems provide a framework for quantifying entropy in finite scenarios by demonstrating that finite-size quantities, like the smoothed min-entropy, can be bounded from below by a sum of asymptotic quantities. Thus, an essential part of a DIQKD security proof reduces to reliably lower bounding the asymptotic quantities in (17).
Beyond DIQKD, (16) is also important for randomness extraction, a process aimed at generating bits that are random and independent of any other system from a quantum experiment [24, 25, 26, 41].
As a concrete example of one-sided randomness extraction, consider the $2$-scenario under the constraint that the underlying, potentially unknown, state $\rho$ violates the CHSH inequality (Clauser-Horne-Shimony-Holt) with a value of $\omega\in[2,2\sqrt{2}]$:
$\langle A_{0}B_{0}+A_{0}B_{1}+A_{1}B_{0}-A_{1}B_{1}\rangle_{\rho}\geq\omega.$ (18)
Notably, in $2$-scenarios, the local polytope is uniquely described by the CHSH inequality alone, which is sufficient to certify non-locality in such settings [31]. Furthermore, in this specific $2$-scenario, with only the CHSH inequality as a constraint, an analytic solution for (16) is achievable in a device-independent context, as demonstrated by [9]. Therefore, we use this as a benchmark in Figure 1.
For the experiment shown in Figure 1 in a $2$-scenario with a CHSH inequality as the constraint, we employ a second-level NPA hierarchy with additional monomials of the form $M_{a|x}N_{b|y}P^{(a)}$. We also apply the speedups for numerical evaluations discussed in Appendix E. This modification yields a valid lower bound that serves as a good approximation. In conclusion, the optimization approach is highly efficient, allowing us to partition the interval into many subintervals while keeping the runtime manageable, even on standard personal computers, with computation times remaining within a few seconds.
III.2 two-sided randomness
In contrast to the one-sided randomness discussed in the previous section, it is also possible to extract randomness from both parties simultaneously. This approach can significantly enhance the power of quantum random number generators. In such generators, where the experiment is performed within a single laboratory, Alice’s and Bob’s laboratories are effectively combined, allowing randomness to be extracted from both devices in parallel. Consequently, this increases the overall efficiency and robustness of the randomness generation process.
In this two-sided extraction scenario, the key quantity of interest is expressed as:
$H(AB|X=\tilde{x},Y=\tilde{y},Q_{E}).$ (19)
This expression represents the conditional entropy of the joint outcomes $A$ and $B$, given fixed measurement settings $X=\tilde{x}$ and $Y=\tilde{y}$, as well as quantum side information $Q_{E}$. This joint entropy quantifies the total randomness that can be extracted from the outcomes of both parties under given conditions, making it particularly relevant for applications that require high degrees of randomness.
As demonstrated in Appendix C, there is a formula for this two-sided randomness extraction problem that closely resembles the formula presented in Theorem II.1. We formalize this relationship by stating it as a corollary.
Corollary III.1.
Lower bounding the conditional von-Neumann entropy
$H(AB|X=\tilde{x},Y=\tilde{y},Q_{E}).$ (20)
can be done by solving the following non-commutative polynomial optimization problem
$\inf\ \frac{1}{\ln 2}$ $(|A|\cdot|B|-1+$ (21)
$\sum_{k=0}^{r}\sum_{a\in A,b\in B}\langle\psi_{Q_{A}Q_{B}Q_{E}}|-\alpha_{k}M_{a|\tilde{x}}N_{b|\tilde{y}}P_{k}^{(ab)}$
$-\beta_{k}P_{k}^{(ab)}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle)$
$\sum_{abxyi}c_{abxyi}\langle\psi_{Q_{A}Q_{B}Q_{E}}|M_{a|x}N_{b|y}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle\geq q_{i}$
$\quad 1\leq i\leq m$
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X}$
$\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X}$
$N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
$[M_{a|x},N_{b|y}]=[M_{a|x},P_{k}^{(a)}]=[N_{b|y},P_{k}^{(a)}]=0$
[…]
$(P_{k}^{(a)})^{2}=P_{k}^{(a)},\quad(P_{k}^{(a)})^{\star}=P_{k}^{(a)}$
$\quad 1\leq k\leq r,\ a\in A.$
Proof.
We consider two applications of III.1, each with distinct constraints and setups.
First, we generalize the CHSH-game presented in subsection III.1 to a two-sided setting. In this case, the setup is governed by the CHSH inequality (18), and we examine a $2$-scenario where both Alice and Bob have two measurement inputs, each yielding one of two possible outcomes. The resulting conditional entropy values are similar to those obtained in [3, Fig. 2] and are visualized in Figure 1. To establish upper bounds in this setting, we formulate an optimization problem as described in [5], enabling us to assess lower bounds on randomness extractable under CHSH constraints.
A more advanced application involves a 2322-scenario constrained by a whole distribution, inspired by the experimental setup in [12], which originally examined only one-sided randomness. Here, we adapt the setup for two-sided randomness extraction. In this scenario, Alice has two mutually unbiased measurement bases, while Bob has one measurement basis that aligns with Alice’s first basis and two additional bases to achieve maximal Bell violation. This configuration allows us to study the randomness generated from both parties simultaneously.
For a concrete example, we consider an honest implementation using an entangled two-qubit state:
$|\Phi\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle).$ (22)
In this setup, Alice and Bob perform measurements in the $x$-$z$ plane with angles specified in Table 1. Each measurement corresponds to a projection operator $P_{\alpha}$, defined by
$P_{\alpha}\coloneqq\frac{1}{2}(\mathds{1}+\sin(\alpha)\sigma_{x}+\cos(\alpha)\sigma_{z}),$ (23)
where $\alpha$ represents the measurement angle and $\sigma_{x}$ and $\sigma_{z}$ are Pauli operators. This experimental configuration provides insights into randomness extraction when multiple measurement bases and entangled states are involved.
Alice Bob
1 00 $\pi/2$
2 $\pi/2$ $\pi/8$
3 - $45\pi/$ 4
Table 1: The tables shows the angles for a protocol for randomness extraction in an honest implementation. Importantly, Alice bases are mutually unbiased.
In the experiment corresponding to Table 1, randomness extraction is done in the following bases
$H(AB|X=0,Y=0,Q_{E}).$ (24)
We show in Figure 1 that we match at least the upper bound with $2$ random bits. Moreover, comparing this result with [3, Fig. 3] questions whether a third basis on Bob’s side has an advantage.
IV Summary
This work focuses on optimizing the conditional von-Neumann entropy within a device-independent framework. From a broader perspective, we ask how much randomness can be extracted from two spatially separated, hypothetical “black boxes” located in Alice’s and Bob’s respective labs, where the outcomes follow the statistical laws of quantum theory. Instead of specifying a concrete quantum experiment represented by a Hilbert space $\mathcal{H}$ and a trace-class operator $\rho$ on $\mathcal{H}$, we rely on the achievable statistics defined by general quantum theory. The core principle underlying the presence of genuine randomness here is the monogamy of correlations, which restricts information sharing in non-signalling theories as demonstrated in [42, 43]. This principle remains applicable even with only limited knowledge of the output statistics in (2). Additionally, bounding the capabilities of Alice, Bob, and Eve to the framework of quantum theory is significant not only for applications but also for advancing our understanding of quantum principles. For instance, [44] exploits the gap between general non-signalling theories and quantum theory, highlighting this difference as a powerful tool.
Our approach serves as a versatile framework for deriving lower bounds on the conditional von-Neumann entropy. The main technical contributions of our work include the application of the integral representation from [27] adapted for this setting, as well as the formulation of a non-commutative polynomial optimization problem, which we solve using the NPA hierarchy [37, 35]. This approach bridges the gap between two established methods: those in [3] and [4]. The technique from [3] has become standard in security proofs for device-independent quantum key distribution (DIQKD) [23] and has since been extended to Petz-Rényi divergences [45]. Compared to [3], our method requires fewer NPA variables, as our projections $P_{k}^{(a)}$ in Theorem II.2 reduce complexity relative to the non-Hermitian operators $Z_{a,i}$ in [3]. This reduction streamlines the operator set in the sum-of-squares cone and enables us to impose additional constraints within the kernel of the map $\varphi$ in [36, Fundamental Lemma]. Moreover, our method builds on discretization techniques for integration from [32], while [3] relies on the Gauss-Radau quadrature method from numerical integration theory. However, the advantages of each integration method for the specific computations required remain open for investigation. The method in [4], though distinct from these approaches, recasts the conditional von-Neumann entropy in (1) as
$H(A|Q_{E})_{\rho_{AQ_{E}}}=S(T_{Q_{A}\to A}(\rho_{Q_{A}Q_{B}}))-S(\rho_{Q_{A}Q_{B}}),$ (25)
interpreting it as an entropy production. Here, the channel $T_{Q_{A}\to A}$ acts as the measurement channel on Alice’s system. The method in [4] establishes lower bounds through a variational approach, employing the Gibbs variational principle and generalized Golden-Thompson inequalities via multivariate trace inequalities [46]. A direct comparison of these conceptually related integral-based methods with the entropy production approach would be a valuable avenue for future research. In section III, we present multiple scenarios to illustrate the effectiveness of our method for randomness extraction. These scenarios include both numerical validations and theoretical comparisons to benchmark results from existing studies.
Firstly, we analyze a $2222$ scenario (two measurement settings per party, each with two outcomes) under CHSH constraints, which is a standard setting in quantum information theory for testing Bell inequalities. We consider this scenario for both one-sided and two-sided randomness extraction, where randomness is certified by outcomes that are either private to one party (one-sided) or to both parties (two-sided). Our results in this $2222$ scenario show strong agreement with the analytical values reported in [9], demonstrating the accuracy and robustness of our approach. Additionally, we explore a more intricate $2322$ scenario, where one party has three measurement settings while the other has two. This scenario is used to examine global randomness extraction, which involves jointly certifying randomness across all outcomes. Our findings, obtained under the statistical framework of Werner states, indicate that this more complex setup does not offer a substantial qualitative advantage over the $2222$ scenario. This suggests that increasing the measurement settings in this configuration may not necessarily enhance randomness extraction, a result that merits further investigation. An intriguing direction for future research would be to systematically compare different configurations - such as the $2222$, $2322$ and $3322$ scenarios (where the latter allows each party three measurement settings with two outcomes) - to determine how these variations impact randomness generation under different assumptions of honest implementation or even a random key basis [30]. This could uncover new insights into the optimal design of randomness extraction protocols. Moreover, it would be valuable to examine randomness extraction in more complex non-local games, such as the magic square game discussed in [44]. These games offer alternative frameworks for randomness certification that might provide unique advantages in terms of security and efficiency in quantum protocols.
V Acknowledgements
GK thanks Mario Berta for very helpful advice around the topic. GK thanks Marco Tomamichel for hosting him at CQT Singapore in fall 2024 and Ian George for many helpful discussions around (DI)QKD. The authors thank Omar Fawzi for a discussion around the topic on an early stage of the project.
Appendix A Proof of Theorem II.1
From the tools developed in [32], we use the following estimate with the definition
$y_{k}:=\sup_{0\leq P\leq\mathds{1}}\operatorname{tr}[P(\sigma t_{k}-\rho)]$ (26)
and then
$\int_{\mu}^{\lambda}\frac{ds}{s}\operatorname{tr}^{+}[\sigma s-\rho]\leq y_{1}[(1+\frac{t_{1}}{t_{2}-t_{1}})\log\frac{t_{2}}{t_{1}}-1]+y_{r}[1-\frac{t_{r-1}}{t_{r}-t_{r-1}}\log\frac{t_{r}}{t_{r-1}}]$ (27)
$\hskip 142.26378pt+\sum_{k=2}^{r-1}y_{k}[(1+\frac{t_{k}}{t_{k+1}-t_{k}})\log\frac{t_{k+1}}{t_{k}}-\frac{t_{k-1}}{t_{k}-t_{k-1}}\log\frac{t_{k}}{t_{k-1}}].$
It is true for all $\mu>0$ and can be summarized with the following coefficients
$\alpha_{k}=\begin{cases}-[(1+\frac{t_{1}}{t_{2}-t_{1}})\log\frac{t_{2}}{t_{1}}-1]&k=1\\ -[1-\frac{t_{r-1}}{t_{r}-t_{r-1}}\log\frac{t_{r}}{t_{r-1}}]&k=r\\ -[(1+\frac{t_{k}}{t_{k+1}-t_{k}})\log\frac{t_{k+1}}{t_{k}}-\frac{t_{k-1}}{t_{k}-t_{k-1}}\log\frac{t_{k}}{t_{k-1}}]&\text{else}.\end{cases}$
and for $\beta_{k}$
$\beta_{k}=\begin{cases}[(1+\frac{t_{1}}{t_{2}-t_{1}})\log\frac{t_{2}}{t_{1}}-1]\cdot t_{1}&k=1\\ [1-\frac{t_{r-1}}{t_{r}-t_{r-1}}\log\frac{t_{r}}{t_{r-1}}]\cdot t_{r}&k=r\\ [(1+\frac{t_{k}}{t_{k+1}-t_{k}})\log\frac{t_{k+1}}{t_{k}}-\frac{t_{k-1}}{t_{k}-t_{k-1}}\log\frac{t_{k}}{t_{k-1}}]\cdot t_{k}&\text{else}.\end{cases}$
For the interval $[0,\mu]$, we use the convexity of the function $s\mapsto\operatorname{tr}^{+}[\sigma s-\rho]$ from [32, Lem. 1] and the fact that for positive operators $\sigma,\rho\in\mathcal{P}(\mathcal{H})$, the function $s\mapsto\operatorname{tr}^{+}[\sigma s-\rho]$ has the limit $\operatorname{tr}^{+}[\sigma s-\rho]\to 0$ as $s\to 0$. Thus, a simple convex interpolation on the interval $[0,\mu]$ gives in addition the following relations: $\alpha_{0}=-1$ and $\beta_{0}=\mu$. Inserting now the definition of $y_{k}$ from (26) into (27) and simplifying everything with the definitions for $\alpha_{k}$ and $\beta_{k}$ for $0\leq k\leq r$ yields the following upper estimate for the relative entropy
$D(\rho\|\sigma)$ $\leq\frac{1}{\ln 2}(\operatorname{tr}[\rho-\sigma]+\sup_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\sum_{k=0}^{r}\operatorname{tr}[P_{k}(\alpha_{k}\rho+\beta_{k}\sigma)]+\ln\lambda+1-\lambda),$ (28)
which is exactly (12) and thus the assertion.
Appendix B Proof Theorem II.2
For simplicity we neglect all the way the whole constraints regadarding Bell-type expressions in (15). Those constraints are well-known and came along many times in similar calculations regarding the device independent bounds on Bell violations. We start with the following expression, which is basically the step in (14), which is certified by Theorem II.1 and $\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}})$
$\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}H(A|X=\tilde{x},Q_{E})_{\rho_{AQ_{E}|\tilde{x}}}=$ $\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}-D(\rho_{AQ_{E}|\tilde{x}}\|\mathds{1}_{A}\otimes\rho_{Q_{E}})$ (29)
$\geq$ $\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\frac{1}{\ln 2}(\operatorname{tr}[\mathds{1}_{A}\otimes\rho_{Q_{E}}-\rho_{AQ_{E}|\tilde{x}}]$
$\quad+\sum_{k=0}^{r}\operatorname{tr}[P_{k}(-\alpha_{k}\rho_{AQ_{E}|\tilde{x}}-\beta_{k}\mathds{1}_{A}\otimes\rho_{Q_{E}})].$
In a next step, we use the fact that
$\operatorname{tr}[\mathds{1}_{A}\otimes\rho_{Q_{E}}-\rho_{AQ_{E}|\tilde{x}}]=|A|-1$ (30)
and that each of the $P_{k}$ is a projector in a classical quantum system such that we can write
$P_{k}=\sum_{a\in A}|a\rangle\langle a|\otimes P_{k}^{(a)}.$ (31)
For some positive operator $P_{k}^{(a)}$. Calculating $(P_{k}^{(a)})^{2}$ then yields that even the $P_{k}^{(a)}$ have to be projections if $P_{k}$ is a projection. In addition we use the following calculation for $\rho_{AQ_{E}|\tilde{x}}$
$\rho_{AQ_{E}|\tilde{x}}$ $=\sum_{a\in A}p(a|x)|a\rangle\langle a|\otimes\rho_{Q_{E}|A=a,\tilde{x}}$ (32)
$=\sum_{a\in A}|a\rangle\langle a|\otimes\operatorname{tr}_{Q_{A}Q_{B}}(\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes\mathds{1}_{B}).$
Before we proceed we briefly argue that of course Eve’s marginal does not depend on $\tilde{x}$. This is the non-signalling property of quantum theory or one may calculate it directly
$\rho_{Q_{E}|\tilde{x}}$ $=\operatorname{tr}_{A}[\rho_{AQ_{E}|\tilde{x}}]$ (33)
$=\operatorname{tr}_{A}[\sum_{a\in A}p(a|x)|a\rangle\langle a|\otimes\rho_{Q_{E}|A=a,\tilde{x}}]$
$=\sum_{a\in A}\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes\mathds{1}_{Q_{B}}]$
$=\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}\sum_{a\in A}M_{a|\tilde{x}}\otimes\mathds{1}_{Q_{B}}]$
$=\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}\mathds{1}_{Q_{A}}\otimes\mathds{1}_{Q_{B}}]$
$=\rho_{Q_{E}}.$
Now we insert (30),(32) and (33) into (29) to get
$\ldots$ $=\frac{1}{\ln 2}\big{(}|A|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{a},\ldots,P_{r}^{a}}\sum_{k=0}^{r}-\alpha_{k}\operatorname{tr}[\sum_{a^{\prime}\in A}|a^{\prime}\rangle\langle a^{\prime}|\otimes\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes\mathds{1}_{Q_{B}Q_{E}}]\sum_{a\in A}|a\rangle\langle a|\otimes P_{k}^{(a)}]$ (34)
$\hskip 170.71652pt+\sum_{k=0}^{r}-\beta_{k}\operatorname{tr}[\mathds{1}_{A}\otimes\rho_{Q_{E}}\sum_{a\in A}|a\rangle\langle a|\otimes P_{k}^{(a)}]\big{)}$
$=\frac{1}{\ln 2}\big{(}|A|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{a},\ldots,P_{r}^{a}}\sum_{k=0}^{r}\sum_{a\in A}-\alpha_{k}\operatorname{tr}[|a\rangle\langle a|\otimes\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes\mathds{1}_{Q_{B}Q_{E}}]P_{k}^{(a)}]$
$\hskip 170.71652pt+\sum_{k=0}^{r}\sum_{a\in A}-\beta_{k}\operatorname{tr}[|a\rangle\langle a|\otimes\rho_{Q_{E}}P_{k}^{(a)}]\big{)}$
$=\frac{1}{\ln 2}\big{(}|A|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{a},\ldots,P_{r}^{a}}\sum_{k=0}^{r}\sum_{a\in A}-\alpha_{k}\operatorname{tr}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes\mathds{1}_{Q_{B}}\otimes P_{k}^{(a)}]-\beta_{k}\operatorname{tr}[\rho_{Q_{A}Q_{B}Q_{E}}\mathds{1}_{Q_{A}}\otimes\mathds{1}_{Q_{B}}\otimes P_{k}^{(a)}]\big{)}.$
Now using the fact that $\rho_{AQ_{E}|\tilde{x}}\leq\mathds{1}_{A}\otimes\rho_{Q_{E}}$ (due to the fact that the state is classical quantum) and thus $\lambda=1$ yields $\ln\lambda+1-\lambda=0$. Moreover, as discussed in Theorem II.1 and the proof in Appendix A, $\mu=0$ can be handled without problems such that the constraints coming from the integral representation in (5) become trivial. Moreover, the operators $P_{k}^{(a)}\in\mathcal{B}(\mathcal{H}_{Q_{E}})$ such that they commute with all local measurements on Alice and Bob’s side respectively. Adding the fact that they are projections from (31), we conclude all the constraints stated in Theorem II.2. Observing the purity argument of $\rho_{Q_{A}Q_{B}Q_{E}}$ by Naimark’s Theorem (e.g. [34, Chap. Dilation Theory]) yields the assertion of the Theorem II.2, if we replace tensor products with commuting operators.
Appendix C Global randomness from the full distribution
We show similarly to Appendix B how to get lower bounds for $H(AB|X=\tilde{x},Y=\tilde{y},Q_{E})$ and abbreviate steps similar to Appendix B. Similarly, we neglect all Bell-type expressions in (15). Applying Theorem II.1 for $\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}})$ yields
$\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}H(AB|X=\tilde{x},Y=\tilde{y}Q_{E})_{\rho_{AQ_{E}|\tilde{x}}}=$ $\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}-D(\rho_{ABQ_{E}|\tilde{x},\tilde{y}}\|\mathds{1}_{AB}\otimes\rho_{Q_{E}})$ (35)
$\geq$ $\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\frac{1}{\ln 2}(\operatorname{tr}[\mathds{1}_{AB}\otimes\rho_{Q_{E}}-\rho_{ABQ_{E}|\tilde{x},\tilde{y}}]$
$\quad+\sum_{k=0}^{r}\operatorname{tr}[P_{k}(-\alpha_{k}\rho_{ABQ_{E}|\tilde{x},\tilde{y}}-\beta_{k}\mathds{1}_{AB}\otimes\rho_{Q_{E}})].$
Using the fact that
$\operatorname{tr}[\mathds{1}_{AB}\otimes\rho_{Q_{E}}-\rho_{ABQ_{E}|\tilde{x},\tilde{y}}]=|A|\cdot|B|-1$ (36)
and that each of the $P_{k}$ is a projector in a classical quantum system, we rewrite
$P_{k}=\sum_{a\in A,b\in B}|a,b\rangle\langle a,b|\otimes P_{k}^{(ab)}.$ (37)
with positive operators $P_{k}^{(ab)}\in\mathcal{P}(\mathcal{H}_{E})$. Furthermore, we can apply the following calculations for $\rho_{ABQ_{E}|\tilde{x},\tilde{y}}$
$\rho_{ABQ_{E}|\tilde{x},\tilde{y}}$ $=\sum_{a\in A,b\in B}p(a,b|x,y)|a,b\rangle\langle a,b|\otimes\rho_{Q_{E}|A=a,B=b,\tilde{x},\tilde{y}}$ (38)
$=\sum_{a\in A,b\in B}|a,b\rangle\langle a,b|\otimes\operatorname{tr}_{Q_{A}Q_{B}}(\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes N_{b|\tilde{y}}).$
Now we insert (36) and (38) into (35) to get
$\ldots$ $=\frac{1}{\ln 2}\big{(}|A|\cdot|B|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{a},\ldots,P_{r}^{a}}\sum_{k=0}^{r}-\alpha_{k}\operatorname{tr}[\sum_{a^{\prime}\in A,b^{\prime}\in B}|a^{\prime},b^{\prime}\rangle\langle a^{\prime},b^{\prime}|\otimes\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes N_{b|\tilde{y}}\otimes\mathds{1}_{Q_{E}}]$ (39)
$\hskip 85.35826pt\sum_{a\in A,b\in B}|a,b\rangle\langle a,b|\otimes P_{k}^{(ab)}]+\sum_{k=0}^{r}-\beta_{k}\operatorname{tr}[\mathds{1}_{AB}\otimes\rho_{Q_{E}}\sum_{a\in A,b\in B}|a,b\rangle\langle a,b|\otimes P_{k}^{(ab)}]\big{)}$
$=\frac{1}{\ln 2}\big{(}|A|\cdot|B|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{ab},\ldots,P_{r}^{ab}}\sum_{k=0}^{r}\sum_{a\in A,b\in B}-\alpha_{k}\operatorname{tr}[|a,b\rangle\langle a,b|\otimes\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes N_{b|\tilde{y}}\otimes\mathds{1}_{Q_{E}}]P_{k}^{(ab)}]$
$\hskip 170.71652pt+\sum_{k=0}^{r}\sum_{a\in A,b\in B}-\beta_{k}\operatorname{tr}[|a,b\rangle\langle a,b|\otimes\rho_{Q_{E}}P_{k}^{(ab)}]\big{)}$
$=\frac{1}{\ln 2}\big{(}|A|\cdot|B|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{ab},\ldots,P_{r}^{ab}}\sum_{k=0}^{r}\sum_{a\in A,b\in B}-\alpha_{k}\operatorname{tr}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes N_{b|\tilde{y}}\otimes P_{k}^{(ab)}]$
$\hskip 113.81102pt-\beta_{k}\operatorname{tr}[\rho_{Q_{A}Q_{B}Q_{E}}\mathds{1}_{Q_{A}Q_{B}}\otimes P_{k}^{(ab)}]\big{)}.$
Now using the fact that $\rho_{ABQ_{E}|\tilde{x},\tilde{y}}\leq\mathds{1}_{AB}\otimes\rho_{Q_{E}}$ (due to the fact that the state is classical quantum) and thus $\lambda=1$ yields $\ln\lambda+1-\lambda=0$. Adding all arguments from Appendix B we deduce from Appendix B a lower bound on two-sided global randomness
$H(AB|X=0,Y=0,Q_{E})$ $\geq\frac{1}{\ln 2}\big{(}|A|+|B|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{a},\ldots,P_{r}^{a}}\sum_{k=0}^{r}\sum_{a\in A,b\in B}-\alpha_{k}\operatorname{tr}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes N_{b|\tilde{y}}\otimes P_{k}^{(ab)}]$ (40)
$\hskip 85.35826pt-\beta_{k}\operatorname{tr}[\rho_{Q_{A}Q_{B}Q_{E}}\mathds{1}_{Q_{A}Q_{B}}\otimes P_{k}^{(ab)}]\big{)}.$ (41)
Appendix D Algebraic Perspective on the NPA hierarchy
The formulation presented in Theorem II.2 uses the language of Hilbert spaces and optimization, establishing a framework based on all possible Hilbert space representations for the physical system under consideration. However, for solving the optimization problem more effectively, it is advantageous to recast the problem within the framework of a $C^{\star}$-algebra. To this end, we define the complex free $\star$-algebra, denoted $\mathcal{F}(\mathcal{G})$, generated by a set of symbols $\mathcal{G}$ as described in [38]. In our case, the generating set $\mathcal{G}$ is given by
[…]
According to [38], we can impose a norm on the algebra $\mathcal{F}(\mathcal{G}|\mathcal{R})$ for a set of relations $\mathcal{R}$. This process allows us, via a completion procedure, to obtain a $C^{\star}$-algebra denoted $C^{\star}(\mathcal{G}|\mathcal{R})$. Lemma 4 in [38] establishes that there is an equivalence between this algebraic formulation and an optimization over its representations.
Within a $C^{\star}$-algebra framework, inner approximations of the positive cone are naturally represented as follows:
$\Sigma_{2}^{(1)}\subset\ldots\subset\Sigma_{2}^{(n)}\subset\Sigma_{2}^{(n+1)}\subset\ldots\subset C^{\star}(\mathcal{G}|\mathcal{R})^{+}.$ (43)
Here, the sequence $\{\Sigma_{2}^{(n)}\}$ represents nested inner approximations of the positive cone $C^{\star}(\mathcal{G}|\mathcal{R})^{+}$. Moving to the dual cones of each $\Sigma_{2}^{(n)}$, we arrive at a corresponding sequence:
$(\Sigma_{2}^{(1)})^{\star}\supset\ldots\supset(\Sigma_{2}^{(n)})^{\star}\supset(\Sigma_{2}^{(n+1)})^{\star}\supset\ldots\supset\mathcal{P}(C^{\star}(\mathcal{G}|\mathcal{R})),$ (44)
where $C^{\star}(\mathcal{G}|\mathcal{R})^{+}$ denotes the natural positive cone of the $C^{\star}$-algebra $C^{\star}(\mathcal{G}|\mathcal{R})$, and $\mathcal{P}(C^{\star}(\mathcal{G}|\mathcal{R}))$ denotes the space of positive, linear, and continuous functionals on $C^{\star}(\mathcal{G}|\mathcal{R})$. By choosing appropriate subsets $\Sigma_{2}^{(n)}$, we can ensure that the dual cone $(\Sigma_{2}^{(n)})^{\star}$ is representable as a semidefinite program (SDP), facilitating a two-step solution approach.
First, we formulate the initial optimization problem as follows:
$c^{\star}\coloneqq\inf$ $\omega(F_{0})$ (45)
s.t. $\omega(F_{i})\leq f_{i}\quad 1\leq i\leq m$
$\omega\in\mathcal{P}(C^{\star}(\mathcal{G}|\mathcal{R})).$
In the first relaxation, we approximate the problem by replacing $\mathcal{P}(C^{\star}(\mathcal{G}|\mathcal{R}))$ with the dual cone $(\Sigma_{2}^{(n)})^{\star}$, yielding
$c^{\star}\geq\inf$ $\omega(F_{0})$ (46)
s.t. $\omega(F_{i})\leq f_{i}\quad 1\leq i\leq m$
$\omega\in(\Sigma_{2}^{(n)})^{\star}.$
In the second step, we exploit the SDP-representability of $(\Sigma_{2}^{(n)})^{\star}$ within a finite-dimensional matrix algebra, constructed via a set of operators $\{K_{j}\}$. By mapping each condition $\{F_{i}\}$ to specific matrices in this SDP framework, we arrive at the final SDP formulation:
$\inf$ $\operatorname{tr}(\rho M_{F_{0}})$ (47)
s.t. $\operatorname{tr}(\rho M_{F_{i}})\geq f_{i}\quad 1\leq i\leq m$
$\operatorname{tr}(\rho K_{j})=0\quad 1\leq j\leq n$
$\rho\geq 0.$
An illustrative example of this methodology is the sum-of-squares (SOS) cones, which were initially introduced in [35, 37]. For SOS cones, the Python package described in [39] automates the translation of generators and relations into an SDP within a matrix algebra framework.
Appendix E Additional discussion for the numerics
We repeat in this section the resulting optimization problem from Theorem II.2
$H(A|X=\tilde{x},Q_{E})\geq\inf$ $\frac{1}{\ln 2}(|A|-1+\sum_{k=0}^{r}\sum_{a\in A}\langle\psi_{Q_{A}Q_{B}Q_{E}}|-\alpha_{k}M_{a|x}P_{k}^{(a)}-\beta_{k}P_{k}^{(a)}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle)$ (48)
$\sum_{abxyi}c_{abxyi}\langle\psi_{Q_{A}Q_{B}Q_{E}}|M_{a|x}N_{b|y}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle\geq q_{i}\quad 1\leq i\leq m$
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X}$
$\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X}$
$N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
[…]
$(P_{k}^{(a)})^{2}=P_{k}^{(a)},\quad(P_{k}^{(a)})^{\star}=P_{k}^{(a)}\quad 1\leq k\leq r,\ a\in A.$
As discussed in Appendix D, the relations
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X}$ (49)
$\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X}$
$N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
[…]
$(P_{k}^{(a)})^{2}=P_{k}^{(a)},\quad(P_{k}^{(a)})^{\star}=P_{k}^{(a)}\quad 1\leq k\leq r,\ a\in A.$
become equality conditions in (47) expressed within the $\{K_{j}\}$. We are now left with the objective function
$\frac{1}{\ln 2}\left(|A|-1+\sum_{k=0}^{r}\sum_{a\in A}\langle\psi_{Q_{A}Q_{B}Q_{E}}|-\alpha_{k}M_{a|x}P_{k}^{(a)}-\beta_{k}P_{k}^{(a)}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle\right),$ (50)
along with some associated constraints. As discussed in [3, Rem. 2.6], we can implement several optimizations to accelerate the numerical evaluations:
  1. 1.
    Modifying the infimum and one of the summations in (50) provides lower bounds, reducing the number of required operators $P_{k}^{(a)}$ to just $\|A\|$. This adjustment significantly accelerates the numerical computations, allowing for the use of a grid with any desired precision.
  2. 2.
    We expand the generating set for the sum-of-squares cone and NPA level $2$ by incorporating additional monomials. In alignment with the approach in [3], we include elements of the form $P_{k}M_{a|x}N_{b|y}$, which enriches the algebraic structure and improves convergence properties.
  3. 3.
    Finally, we can set the NPA matrices to be real-valued, which simplifies computations further.
Applying these optimizations enables all computations to complete within seconds.