### v2

# Reliable Entropy Estimation from Observed Statistics for Device-Independent Quantum Cryptography

Abstract
We introduce a numerical framework for reliably estimating conditional von Neumann entropies, a central quantity in device-independent quantum cryptography and randomness generation. Our method is based on semidefinite relaxations derived from the Navascués–Pironio–Acín hierarchy and provides entropy bounds directly from observed statistics, assuming only the validity of quantum mechanics. The approach builds on a recent integral representation of entropy and substantially improves efficiency: it requires about half as many support variables as existing methods, while allowing these variables to be chosen projectively, leading to a significant reduction in runtime. These advances enable the practical certification of randomness and security even in noisy conditions, while integrating seamlessly with modern entropy accumulation theorems. Consequently, our framework becomes a versatile tool for quantum cryptographic protocols, broadening the possibilities for secure communication in untrusted environments.
I Introduction
Almost a century ago, Max Born [1] famously suggested that the unpredictability observed in individual quantum measurements cannot always be explained by an incomplete knowledge about the experimental setup alone. Instead, this unpredictability can be an intrinsic property of nature itself—that is, true randomness exists. This idea is strikingly demonstrated by a successful Bell test.
Regarded from the perspective of quantum mechanics, a maximal violation of a Bell inequality like CHSH [2] demonstrates two key points. On one hand, it rules out the existence of hidden variables that could secretly determine a outcome. On the other, it shows that the correlated yet individually unpredictable measurement outcomes—emerging from a pure, maximally entangled state—are truly random and uncorrelated with any third party not involved in the experiment.
Today, Bell-type correlation experiments that are capable of closing a wide range of loopholes have moved beyond mere thought experiments—they are now a practical reality (see [3, 4, 5, 6, 7, 8] and references therein). However, since all physical experiments must contend with noise and decoherence, a maximal violation indicative of a pure state will most likely remain an idealized model. This necessitates the task of characterizing and certifying the amount of randomness present in observed noisy real-world data—an issue we address in this work.
Recognizing that the outcomes of a quantum measurement can be genuinely random marked not only a fundamental shift in our understanding of the world but also, much later, paved the way for the development of numerous information-theoretic protocols that would be impossible to achieve under the framework of any classical theory [9]. Most prominently, the field of quantum cryptography evolved from this, which by today ranges from the development of near-commercial technologies to foundational research pushing the limits of current capabilities.
Figure 1: Basic setting of a quantum mechanical correlation experiment for the generation of randomness or secret key. A source distributes a quantum state $\rho_{Q_{A}Q_{B}Q_{E}}$ between two legit users Alice and Bob, and an adversary Eve. Depending on inputs $x$ and $y$, Alice and Bob perform measurements resulting in classical random variables AA and $B$ to which Eve holds quantum side information. Our main focus is to estimate the securely asymptotically extractable randomness of AA and/or $B$ based on only considering the observed probability distributions $p(a,b|x,y)$.
Recent advancements in quantum cryptography, particularly in device-independent quantum key distribution (DIQKD) and randomness extraction (see e.g. [10, 11] and references therein), harness the intrinsic unpredictability of quantum systems to secure communication against eavesdroppers. DIQKD protocols rely on the principle that correlations observed in measurement outcomes from entangled quantum systems cannot be explained by any local hidden variable theory, as shown by violations of Bell inequalities.
This fundamental characteristic implies that the conditional distribution
$p(a,b|x,y),$ (1)
see Fig. 1, cannot be reproduced if $a$ and $b$ were deterministic outputs available to an eavesdropper for each input pair $(x,y)$, with values governed by a shared random variable, such as a shared coin toss. Instead, quantum non-locality—revealed through such correlations—ensures that outcomes cannot be explained by classical models. This establishes true randomness as a consequence of quantum theory. The DIQKD approach thus offers an additional layer of security by requiring no trust in the measurement devices themselves. Given this framework, device-independent protocols make the accurate bounding of entropy under non-asymptotic conditions a central challenge [12]. The need to rigorously quantify the randomness of measurement outcomes—especially in real-world scenarios where noise and finite sample sizes introduce deviations—underscores the importance of non-asymptotic entropy bounds. Such bounds provide critical insights into the level of security achievable in device-independent settings, where only the observed distribution and the validity of quantum theory are assumed [13, 14] (of course in an actual experiment many more assumptions on the devices come into play; we restrict here on the mathematical framework). As research progresses, these entropy bounds are becoming essential for verifying and implementing DIQKD protocols in practical cryptographic systems, thereby expanding the possibilities for secure communication in untrusted environments [15, 16, 17].
In this work, we contribute to the field of device-independent cryptography by introducing a numerical method that provides provable lower bounds on the conditional von Neumann entropy, which quantifies the asymptotic amount of usable randomness extractable from a given experiment. Specifically, we bound the entropy
$H(A|Q_{E})_{\rho_{AQ_{E}}}\coloneqq S(\rho_{AQ_{E}})-S(\rho_{Q_{E}}),$ (2)
where $A$ is a classical random variable held by Alice and $Q_{E}$ is a quantum system held by Eve. Our method enables certification of randomness using only the observed measurement outcomes from Alice and Bob, under the sole assumption that the experiment is governed by quantum theory. The core idea of our method is to transform a novel class of integral representations of the relative entropy [18] into a sequence of non-commutative optimization problems. These problems can be approximated using the Navascués-Pironio-Acín (NPA) hierarchy of semidefinite programs [19, 20]. While the approximation is still computationally demanding for larger systems, our method offers a significant improvement in efficiency compared to previous approaches.
Computing conditional von Neumann entropies in the device independent setting has been an outstanding open problem for quite some time [12]. Only recently the first methods for this problem were developed by Tan et al. [21] and by Brown, Fawzi, and Fawzi [22], shortly after. Both are suffering from serious performance issues when it comes to analyse protocols with more than an minimal amount of inputs and outputs on each side. This limitation is due the fact that both methods require the simultaneous analysis of many non-commutative variables and/or high-degree polynomials. We have the capability to fix this. In its most economical version, our method can give direct estimates on the von Neumann entropy, whilst requiring minimal computational resources. In detail we need the same amount of resources as in the NPA-based computation of min-entropies [23], which is a lower but usually suboptimal bound to the von Neumann entropy. Nevertheless, if more resources can be allocated we can enhance the approximation quality and convergence speed to the optimum.
In general, the demand for such numerical tools has increased recently evermore due to significant progress in entropy accumulation theorems [13, 14], which provide concrete solutions for randomness extraction in finite-size regimes by reducing it to the asymptotic quantity in (2). Thus, after applying an entropy accumulation theorem, the remaining task is to calculate provable lower bounds for the conditional von-Neumann entropy (2) for specific noise models and input-output data scenarios. Furthermore, by confining the analysis to projective operators, our method achieves a high degree of computational efficiency, making it suitable for real-world applications.
This paper is structured as follows: In section Sec. II, we will present the underlying mathematical framework for our method. We will introduce and analyse an integral representation of the underlying optimization problem and introduce approximations, that are linear in the underlying state, see Thm. II.1. This enables us to formulate the asymptotically extractable randomness as a non-commutative polynomial optimization problem, Thm.II.2. The following sections are dedicated to applications. In Sec.III we will revisit the well known 2222-setting. Here, Alice has 2 measurements with 2 outcomes, and Bob has 2 measurements with 2 outcomes. In section Sec. IV we will focus on settings with more measurements and outcomes. In detail, we will compute the asymptotically extractable randomness for settings in which the CGLMP inequality and the I3322 are measured. For these cases, previous numerical methods do not longer perform well [24]. Our method however does, and we demonstrate that direct bounds on the von Neumann entropy certify much higher randomness extraction rates than a comparable ansatz based on the min-entropy. As a last demonstration, presented in section Sec. V we will test our method on real world experimental data. The data comes from a test run of the experiment [7], which firstly demonstrated the feasibility of DI-QKD with distant users. We show that using the statistics of more than 2 measurements can actually increase the amount of certifiable randomness when compared to a simple test of the CHSH inequality [23].
II Bounding the von-Neumann entropy
In the following we denote by $\mathcal{B}(\mathcal{H})$ the set of (bounded) linear operators on a Hilbert space $\mathcal{H}$ and $\mathcal{S}(\mathcal{H})$ the set of quantum states on $\mathcal{H}$, i.e. all positive operators with unit trace. The trace on $\mathcal{B}(\mathcal{H})$ is denoted as $\operatorname{tr}[\cdot]$. The positive semidefinite cone is denoted by $\mathcal{P}(\mathcal{H})$. Moreover, any self adjoint operator $A\in\mathcal{B}(\mathcal{H})$, can be uniquely decomposed as a difference $A=A^{+}-A^{-}$ of Hilbert-Schmidt orthogonal positive operators $A^{+}$ and $A^{-}$. Let $\operatorname{tr}^{+}[A]\coloneqq\operatorname{tr}[A^{+}]$ denote the trace of the positive part of $A$ (similarly $\operatorname{tr}^{-}[A]\coloneqq\operatorname{tr}[A^{-}]=\operatorname{tr}^{+}[-A]$). Note that this is an SDP given by
$\operatorname{tr}^{+}[A]=\sup$ $\operatorname{tr}[PA]$ (3)
$\operatorname{s.t.}$ $0\leq P\leq\mathds{1}.$
[…]
by P. E. Frenkel [18]. For two positive operators $\sigma,\rho\in\mathcal{P}(\mathcal{H})$ acting on a Hilbert space $\mathcal{H}$ the r.h.s. of (4) can be expressed by the integration
$D(\rho\|\sigma)=\frac{1}{\ln 2}\big{(}\operatorname{tr}[\rho-\sigma]+\int_{-\infty}^{+\infty}dt\frac{\operatorname{tr}^{-}[(1-t)\rho+t\sigma]}{|t|(t-1)^{2}}\big{)}.$ (5)
It was shown by Jenčová [25], that the formulation (5) also holds for infinite-dimensional yet separable Hilbert spaces. It can be further rewritten in a form that is more convenient for our purposes (see A.1)
$D(\rho\|\sigma)$ $=\frac{1}{\ln 2}\big{(}\operatorname{tr}[\rho-\sigma]+\int_{\mu}^{\lambda}\frac{ds}{s}\operatorname{tr}^{+}[\sigma s-\rho]$ (6)
$\hskip 28.45274pt+\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma]\big{)}.$
Here, the boundaries of the integration range $\mu,\lambda\in\mathbb{R}_{\geq 0}$ have to obey the operator valued inequality
$\mu\sigma\leq\rho\leq\lambda\sigma.$ (7)
In the DI settings we want to analyse, the states $\rho$ and $\sigma$ of interest arise from the protocol under consideration, here a generic choice for valid $\lambda$ and $\mu$ can be given.
For the remainder of this work, we assume tripartite experiments involving Alice, Bob, and Eve, modeled by the axioms of quantum theory as a tripartite state expressed by a positive operator on a Hilbert space with trace equal to $1$,
$\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{QE}).$ (8)
Importantly, the argument for the device independence specifies nothing concrete about the Hilbert space $\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{QE}$. In the following we follow the example of generating a raw-key in a DIQKD protocol, even though our methods can be applied for a larger class of problems in section III. Afterwards, we briefly discuss some simple generalizations.
In section III we elaborate on this assumption for certain QKD scenarios. In a DIQKD protocol we usually have some measurements $\mathcal{X}$ on $\mathcal{S}(\mathcal{H}_{Q_{A}})$ with finite outcome set $A$ and similarly $\mathcal{Y}$ on $\mathcal{S}(\mathcal{H}_{Q_{B}})$ with finite outcome set $B$ respectively. Measurements, see Fig.1, are generally understood as sets of POVMs $\{M_{a|x}\}_{a\in A}$ for each $x\in\mathcal{X}$ and similarly $\{M_{b|y}\}_{b\in B}$ for each $y\in\mathcal{Y}$. From a more abstract perspective a measurement could also be understood as channel that maps from a commutative algebra into the observable algebra of a quantum system. In the most simplest form of a protocol, the raw-key is now constructed by one specific measurement $\tilde{x}\in\mathcal{X}$. This can be generalized later on [26], but we skip the technical details of this for now. The asymptotically extractable randomness of the fixed measurement $\tilde{x}\in\mathcal{X}$ can be described as [^2] For simplicity we abbreviate $\rho_{AQ_{E}|\tilde{x}}\equiv\rho_{AQ_{E}|X=\tilde{x}}$
$H(A|X=\tilde{x},Q_{E})_{\rho_{AQ_{E}|\tilde{x}}}=-D(\rho_{AQ_{E}|\tilde{x}}\|\mathds{1}_{A}\otimes\rho_{Q_{E}}).$ (9)
Hence, we have to lower bound $D(\rho_{AQ_{E}|\tilde{x}}\|\mathds{1}_{A}\otimes\rho_{Q_{E}})$. Specifically, device independent lower bounds satisfy the following prototypical type of optimization problem
$\inf$ $-D(\rho_{AQ_{E}|\tilde{x}}\|\mathds{1}_{A}\otimes\rho_{Q_{E}})$ (10)
$\sum_{abxyi}c_{abxyi}\operatorname{tr}[M_{a|x}\otimes N_{b|y}\rho_{Q_{A}Q_{B}}]\geq q_{i},\quad 1\leq i\leq m$
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X}$
$\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X}$
$N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
$\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}})$
whereby we do not aim to specify the concrete underlying Hilbert space. Moreover, the second line can be understood as information about the joint statistics of Alice and Bob, which is nothing more than (1). A simple but sufficient and common example would be that the $c_{abxyi}\in\mathbb{R}$ coefficients form a Bell inequality (see e.g. [27]). In the following we need to estimate the program (10) with an actually quantifiable program. For this purpose, we use the integral representation (6) and derive an upper bound for the relative entropy (4) with an operator valued optimization problem.
Theorem II.1 (Upper bounds for relative entropy).
Consider two positive operators $\rho,\sigma\in\mathcal{P}(\mathcal{H})$ such that
$\mu\sigma\leq\rho\leq\lambda\sigma$ (11)
for arbitrary $\mu,\lambda\in\mathbb{R}_{\geq 0}$. Then we can upper bound
$\int_{0}^{\lambda}\frac{ds}{s}\operatorname{tr}^{+}[\sigma s-\rho]\leq\sup_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\sum_{k=0}^{r}\operatorname{tr}[P_{k}(\alpha_{k}\rho+\beta_{k}\sigma)]$ (12)
for sets of real numbers $\alpha_{0},\ldots,\alpha_{r},\beta_{0},\ldots,\beta_{r}\in\mathbb{R}$, which can be easily precomputed.
In particular the estimate (12) yields
$D(\rho\|\sigma)$ $\leq\frac{1}{\ln 2}(\operatorname{tr}[\rho-\sigma]$ (13)
$+\sup_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\sum_{k=0}^{r}\operatorname{tr}[P_{k}(\alpha_{k}\rho+\beta_{k}\sigma)]$
$+\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma]).$
Proof.
See Appendix A. ∎
The Theorem II.1 provides a sequence of upper bounds that converge to the relative entropy. Specifically, this sequence of optimization problems can be understood as an application of [28]. We emphasize that a finite $\lambda$ is, at least in finite dimensions, equivalent to the statement that the relative entropy is finite. This is a crucial component in these optimization problems, allowing for effective numerical analysis. In the case of conditional von-Neumann entropy, a finite $\lambda$ always exists, as for any bipartite state $\rho_{AB}$ we have the fundamental operator inequality (see, e.g., [29, Appendix A]):
$\rho_{AB}\leq d_{A}\mathds{1}_{A}\otimes\rho_{B}.$ (14)
Now, let us examine the specific optimization problem in (10). It can be lower-bounded using Theorem II.1 as an optimization over operators $P_{0},\ldots,P_{r}$. In particular, we observe that the optimal operators $P_{k}$, $0\leq k\leq r$, are projections, as they correspond to projections onto the positive part in the Jordan decomposition of the Hermitian operator $\alpha_{k}\rho+\beta_{k}\sigma$. Furthermore, by substituting the specific operators $\rho=\mathds{1}_{A}\otimes\rho_{Q_{E}}$ and $\sigma=\rho_{AQ_{E}|\tilde{x}}$, we find that $\lambda=1$ and $\mu$ can be set to 0 without loss of generality (see Appendix A for details). Thus, we can rewrite (10) as
$\inf$ $\inf_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\frac{1}{\ln 2}(\operatorname{tr}[\mathds{1}_{A}\otimes\rho_{Q_{E}}-\rho_{AQ_{E}|\tilde{x}}$] (15)
$+$ $\sum_{k=0}^{r}\operatorname{tr}[P_{k}(-\alpha_{k}\rho_{AQ_{E}|\tilde{x}}-\beta_{k}\mathds{1}_{A}\otimes\rho_{Q_{E}})]$
$\sum_{abxyi}c_{abxyi}\operatorname{tr}[M_{a|x}\otimes N_{b|y}\rho_{Q_{A}Q_{B}}]\geq q_{i},\quad 1\leq i\leq m$
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X},\quad\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X},\quad N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
$\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}})$
$P_{k}^{2}=P_{k}\in\mathcal{P}(\mathcal{H}_{A}\otimes\mathcal{H}_{Q_{E}}).$
An application of Naimark’s dilation theorem additionally allows us to rewrite the state $\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}})$ without loss of generality as a pure state $\psi_{Q_{A}Q_{B}Q_{E}}\in\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}}$, with all measurements also being projective. By carefully examining the classical-quantum structure of $\rho_{AQ_{E}|\tilde{x}}$, the post-measurement state in our specific key basis $\tilde{x}\in\mathcal{X}$ can then be used to reformulate (15) as a non-commutative polynomial optimization in Theorem II.2.
Theorem II.2 (NPO Formulation).
Calculating lower bounds for the conditional von-Neumann entropy (2) in the device independent setting with knowledge about statistics between Alice and Bobs systems can be reformulated into the following non-commutative polynomial optimization problem
$H(A|$ $X=\tilde{x},Q_{E})\geq$ (16)
$\inf$ $\frac{1}{\ln 2}(|A|-1+$
$\sum_{k=0}^{r}\sum_{a\in A}\langle\psi_{Q_{A}Q_{B}Q_{E}}|-\alpha_{k}M_{a|x}P_{k}^{(a)}$
$-\beta_{k}P_{k}^{(a)}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle)$
$\sum_{abxyi}c_{abxyi}\langle\psi_{Q_{A}Q_{B}Q_{E}}|M_{a|x}N_{b|y}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle\geq q_{i}$
$\quad 1\leq i\leq m$
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X},\quad\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X},\quad N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
$[M_{a|x},N_{b|y}]=[M_{a|x},P_{k}^{(a)}]=[N_{b|y},P_{k}^{(a)}]=0$
$\quad b\in B,\ a\in A,\ x\in\mathcal{X},\ y\in\mathcal{Y},0\leq k\leq r$
$(P_{k}^{(a)})^{2}=P_{k}^{(a)},\quad(P_{k}^{(a)})^{\star}=P_{k}^{(a)}$
$\quad 1\leq k\leq r,\ a\in A.$
Proof.
See Appendix B
The polynomial optimization problem (16) can be interpreted as an optimization over the algebraic constraints within it. In a finite-dimensional Hilbert space setting, these constraints can be viewed as generating rules of a $\star$-subalgebra of $\mathcal{B}(\mathcal{H})$, where $\mathcal{H}$ is a Hilbert space (in infinite dimensions, caution is required to handle the topological closure of the algebra, though this remains feasible). Assuming that these relations form an algebra $\mathcal{A}$, we employ the Navascues-Pironio-Acin (NPA) hierarchy [19], to compute outer approximations to the optimization problem (16), thus lower bounds.
The approach involves relaxing the positivity requirement for linear, continuous functionals on $\mathcal{A}$. Rather than requiring positivity across the positive cone $\mathcal{A}_{+}$, we restrict it to a sum-of-squares cone $\Sigma_{2}^{(n)}\subset\mathcal{A}_{+}$ at level $n$. Optimization over this relaxed SDP can then be viewed as optimizing the dual cone $(\Sigma_{2}^{(n)})^{\star}$, which is representable in SDP form and provides a convergent hierarchy of outer approximations to (16). For further details, we refer to [19, 30, 20] and Appendix D. In our applications, we use the software package from [31] and code is available here.
0.750.750.80.80.850.8500.50.511CHSH valueconditional entropylower boundanalytic value
0.750.750.80.80.850.8500.50.5111.51.5CHSH valueconditional entropylower boundupper bound [32]
0.80.80.90.91101122white noiseconditional entropylower bound 2322
Figure 2: The plots illustrate examples of lower bounds for conditional entropies across different scenarios. In the leftmost plot, we present a proof-of-principle example for lower bounds on the conditional entropy $H(A|X=0,Q_{E})_{\rho}$, depending on various values of the CHSH violation in a 2222-scenario (where each of two parties can choose between two measurements with two outcomes). This example is particularly significant because it allows for a verification of the general correctness of the method; as shown in [11], there are analytical values available for this scenario, enabling comparisons with computed results to validate accuracy. The middle plot displays the lower bounds on global randomness, specifically the conditional entropy $H(AB|X=0,Y=0,Q_{E})$, again in a 2222-scenario and upper bounds developed by [32]. Here, the dependency is on the CHSH inequality, and this plot serves to examine the impact of the CHSH violation on the joint conditional entropy of outcomes $A$ and $B$ given measurement settings $X=0$ and $Y=0$. Finally, the rightmost plot demonstrates the quantity $H(AB|X=0,Y=0,Q_{E})$ in a 2322-scenario, where each party has two possible outcomes and Bob has three measurement choices. In this case, we condition on the full distribution of a Werner state depending on dephasing noise. All plots can be generated within seconds to a few minutes of computation on a machine equipped with 12 threads and 128GB of RAM.
III Applications and numerical examples
In this section, we demonstrate the methods and tools developed in section II. To structure this discussion, we divide section III into two parts:
In the first part, we examine a CHSH game scenario with two possible inputs and two possible outputs. Our objective here is to compute $H(A|X=0,Q_{E})$, which is a quantity of particular interest in two main areas: one-sided device-independent randomness extraction (DIRE) and device-independent quantum key distribution (DIQKD). In one-sided DIRE, $H(A|X=0,Q_{E})$ provides a measure of the randomness that can be reliably extracted from Alice’s outcomes, assuming a fixed input $X=0$ and some quantum side information $Q_{E}$. Similarly, in DIQKD, this conditional entropy helps in determining the asymptotic amount of secure randomness extractable from the quantum correlations between Alice and an adversary.
In the second part, we extend this investigation to the two-sided DIRE task, again within a CHSH game framework. Here, the focus shifts to the quantity $H(AB|X=0,Y=0,Q_{E})$, where the conditional entropy now includes both Alice’s and Bob’s outcomes. This joint entropy becomes essential for evaluating the global randomness accessible when both parties contribute to the measurement outcomes in the context of device-independent randomness extraction.
For clarity, we abbreviate scenarios based on the number of inputs and outputs per party. Specifically, a scenario with $n_{a}$ inputs for Alice, $n_{b}$ inputs for Bob, $o_{a}$ possible outcomes for each input of Alice, and $o_{b}$ possible outcomes for each input of Bob is denoted by $n_{a}n_{b}o_{a}o_{b}$. For instance, the CHSH scenario with two inputs and two outputs per party is a 2222-scenario. Furthermore we observe that all quantities under consideration in this work are usable for spot-checking protocols. An interesting direction for future research is to extend this to averaging the entropy terms without spot-checking.
III.1 one-sided randomness
In device-independent quantum information tasks, one-sided (device-independent) randomness is quantified as follows:
$H(A|X=\tilde{x},Q_{E})_{\rho_{AQ_{E}|\tilde{x}}}$, (17)
where this quantity captures the uncertainty in the outcome $A$ when measurement $X$ takes a fixed value $\tilde{x}$, conditioned on quantum side information $Q_{E}$ held by an eavesdropper. This concept is fundamental in the context of device-independent quantum key distribution (DIQKD) and one-sided randomness extraction [32, 33].
In quantum key distribution (QKD), the asymptotic key rate $r_{\rightarrow}$ can be bounded using the Devetak-Winter formula, which states:
$r_{\rightarrow}\geq H(A|E)_{\rho_{AQ_{E}}}-H(A|B)_{\rho_{AB}},$ (18)
where $H(A|E)_{\rho}$ denotes the conditional entropy of $A$ given the eavesdropper’s knowledge $E$, and $H(A|B)_{\rho}$ denotes the conditional entropy of $A$ given the legitimate party $B$. In DIQKD experiments, estimating $H(A|E)_{\rho}$ relies on observed statistics, such as violations of Bell inequalities or even the full probability distribution gathered from the experiment.
This quantity remains significant even in finite-sample-size regimes due to recent advancements in entropy accumulation theorems (EAT) [13, 14]. These theorems provide a framework for quantifying entropy in finite scenarios by demonstrating that finite-size quantities, like the smoothed min-entropy, can be bounded from below by a sum of asymptotic quantities. Thus, an essential part of a DIQKD security proof reduces to reliably lower bounding the asymptotic quantities in (18).
Beyond DIQKD, (17) is also important for randomness extraction, a process aimed at generating bits that are random and independent of any other system from a quantum experiment [15, 16, 17].
As a concrete example of one-sided randomness extraction, consider the 2222-scenario under the constraint that the underlying, potentially unknown, state $\rho$ violates the CHSH inequality (Clauser-Horne-Shimony-Holt) with a value of $\omega\in[2,2\sqrt{2}]$:
$\langle A_{0}B_{0}+A_{0}B_{1}+A_{1}B_{0}-A_{1}B_{1}\rangle_{\rho}\geq\omega.$ (19)
Notably, in 2222-scenarios, the local polytope is uniquely described by the CHSH inequality alone, which is sufficient to certify non-locality in such settings [27]. Furthermore, in this specific 2222-scenario, with only the CHSH inequality as a constraint, an analytic solution for (17) is achievable in a device-independent context, as demonstrated by [34]. Therefore, we use this as a benchmark in Figure 2.
For the experiment shown in Figure 2 in a 2222-scenario with a CHSH inequality as the constraint, we employ a second-level NPA hierarchy with additional monomials of the form $M_{a|x}N_{b|y}P^{(a)}$. We also apply the speedups for numerical evaluations discussed in Appendix E. This modification yields a valid lower bound that serves as a good approximation. In conclusion, the optimization approach is highly efficient, allowing us to partition the interval into many subintervals while keeping the runtime manageable, even on standard personal computers, with computation times remaining within a few seconds.
III.2 two-sided randomness
In contrast to the one-sided randomness discussed in the previous section, it is also possible to extract randomness from both parties simultaneously. This approach can significantly enhance the power of quantum random number generators. In such generators, where the experiment is performed within a single laboratory, Alice’s and Bob’s laboratories are effectively combined, allowing randomness to be extracted from both devices in parallel. Consequently, this increases the overall efficiency and robustness of the randomness generation process.
In this two-sided extraction scenario, the key quantity of interest is expressed as:
$H(AB|X=\tilde{x},Y=\tilde{y},Q_{E}).$ (20)
This expression represents the conditional entropy of the joint outcomes $A$ and $B$, given fixed measurement settings $X=\tilde{x}$ and $Y=\tilde{y}$, as well as quantum side information $Q_{E}$. This joint entropy quantifies the total randomness that can be extracted from the outcomes of both parties under given conditions, making it particularly relevant for applications such as random number generator.
As demonstrated in Appendix C, there is a formula for this two-sided randomness extraction problem that closely resembles the formula presented in Theorem II.1. We formalize this relationship by stating it as a corollary.
Corollary III.1.
Lower bounding the conditional von-Neumann entropy
$H(AB|X=\tilde{x},Y=\tilde{y},Q_{E})_{\rho_{ABQ_{E}|\tilde{x},\tilde{y}}}$
can be done by solving the following non-commutative polynomial optimization problem
$\inf\ \frac{1}{\ln 2}$ $(|A|\cdot|B|-1+$ (21)
$\sum_{k=0}^{r}\sum_{a\in A,b\in B}\langle\psi_{Q_{A}Q_{B}Q_{E}}|-\alpha_{k}M_{a|\tilde{x}}N_{b|\tilde{y}}P_{k}^{(ab)}$
$-\beta_{k}P_{k}^{(ab)}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle)$
$\sum_{abxyi}c_{abxyi}\langle\psi_{Q_{A}Q_{B}Q_{E}}|M_{a|x}N_{b|y}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle\geq q_{i}$
$\quad 1\leq i\leq m$
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X}$
$\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X}$
$N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
$[M_{a|x},N_{b|y}]=[M_{a|x},P_{k}^{(a)}]=[N_{b|y},P_{k}^{(a)}]=0$
$\quad b\in B,\ a\in A,\ x\in\mathcal{X},\ y\in\mathcal{Y},0\leq k\leq r$
$(P_{k}^{(a)})^{2}=P_{k}^{(a)},\quad(P_{k}^{(a)})^{\star}=P_{k}^{(a)}$
$\quad 1\leq k\leq r,\ a\in A.$
Proof.
We consider two applications of III.1, each with distinct constraints and setups.
First, we generalize the CHSH-game presented in subsection III.1 to a two-sided setting. In this case, the setup is governed by the CHSH inequality (19), and we examine a 2222-scenario where both Alice and Bob have two measurement inputs, each yielding one of two possible outcomes. The resulting conditional entropy values are similar to those obtained in [22, Fig. 2] and are visualized in Figure 2. To establish upper bounds in this setting, we formulate an optimization problem as described in [32], enabling us to assess lower bounds on randomness asymptotically extractable under CHSH constraints.
A more advanced application involves a 2322-scenario constrained by a whole distribution, inspired by the experimental setup in [7], which originally examined only one-sided randomness. Here, we adapt the setup for two-sided randomness extraction. In this scenario, Alice has two mutually unbiased measurement bases, while Bob has one measurement basis that aligns with Alice’s first basis and two additional bases to achieve maximal Bell violation. This configuration allows us to study the randomness generated from both parties simultaneously.
For a concrete example, we consider an honest implementation using an entangled two-qubit state:
$|\Phi\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle).$ (22)
In this setup, Alice and Bob perform measurements in the $x$-$z$ plane with angles specified in Table 1. Each measurement corresponds to a projection operator $P_{\alpha}$, defined by
$P_{\alpha}\coloneqq\frac{1}{2}(\mathds{1}+\sin(\alpha)\sigma_{x}+\cos(\alpha)\sigma_{z}),$ (23)
where $\alpha$ represents the measurement angle and $\sigma_{x}$ and $\sigma_{z}$ are Pauli operators. This experimental configuration provides insights into randomness extraction when multiple measurement bases and entangled states are involved.
Alice Bob
1 0 $\pi/2$
2 $\pi/2$ $\pi/8$
3 - $5\pi/4$
Table 1: The tables shows the angles for a protocol for randomness extraction in an honest implementation. Importantly, Alice bases are mutually unbiased.
In the experiment corresponding to Table 1, randomness extraction is done in the following bases
$H(AB|X=0,Y=0,Q_{E}).$ (24)
We show in Figure 2 that we match at least the upper bound with $2$ random bits. Moreover, comparing this result with [22, Fig. 3] questions whether a third basis on Bob’s side has an advantage.
IV Beyond CHSH
CHSH has long been the primary Bell inequality used for device-independent randomness. This is mainly because the $2222$-scenarios enjoy the powerful advantage that the $C^{\star}$-algebra of two projections is well-understood and is essentially as intricate as qubit-based information theory. Many seminal results have been obtained by reducing the device-independent scenario on the algebra of two projections back to the qubit case [11].
However, one can easily conceive of protocols featuring significantly different arrangements of input-output scenarios, which then naturally give rise to different families of Bell inequalities whenever quantum nonlocality arises. In such cases, a straightforward reduction to qubits or even qudits is often not feasible, necessitating either the use of self-testing methods for specific distributions, states and measurements or the acceptance of the full generality of solving the optimization in Theorem II.2.
In a recent result [24], DIQKD protocols beyond qubits were investigated, proposing key rates for more general scenarios than the CHSH and $2222$-cases. However, computing bounds on these key rates using the Brown–Fawzi–Fawzi (BFF) method can be challenging with standard computational resources. For this reason, the results in [24] were obtained by evaluating the min-entropy (i.e., by optimizing the guessing probability). With stronger computational resources, we present a comparison of our method with BFF for I3322 in Figure 4, and with BFF for CGLMP in Figure 6 and show that our method is on a comparable level faster and gives better bounds. Since our focus here is on the fundamental theoretical building blocks of entropy estimation in device-independent scenarios, we concentrate on calculating lower bounds on the asymptotically extractable randomness but in terms of bounds which are converging towards the conditional von Neumann entropy. As we will show numerically in the following, our tools are able to estimate the randomness for the CGLMP 2233-Bell inequality [35] and the I3322 Bell-inequality [36] significantly beyond the bounds from the min-entropy.
$p_{\operatorname{guess}}$ $(A|E)=$ (25)
$\sup$ $\sum_{a\in A}\langle\psi_{Q_{A}Q_{B}Q_{E}}|M_{a|\tilde{x}}C_{a}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle$
$\sum_{abxyi}c_{abxyi}\langle\psi_{Q_{A}Q_{B}Q_{E}}|M_{a|x}N_{b|y}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle\geq q_{i}$
$\quad 1\leq i\leq m$
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X}$
$\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X}$
$N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
$[M_{a|x},N_{b|y}]=[M_{a|x},C_{a^{\prime}}]=[N_{b|y},C_{a^{\prime}}]=0$
$\quad b\in B,\ a,a^{\prime}\in A,\ x\in\mathcal{X},\ y\in\mathcal{Y}$
$\sum_{a}C_{a}=\mathds{1},\ C_{a}\geq 0,\quad a\in A$
and the min entropy is then defined as
$H_{\operatorname{min}}(A|E)\coloneqq-\log p_{\operatorname{guess}}(A|E).$ (26)
IV.1 Randomness from CGLMP
Our first example concerns the CGLMP Bell inequality for a $2233$ scenario [35, eqs. (4) and (5)]. This is a Bell inequality for two parties, each having two measurement settings and three outcomes per setting. If we define, for generally correlated probability distributions, the expression
$P(A_{a}=B_{b}+k)\coloneqq\sum_{j=0}^{d-1}P\bigl{(}A_{a}=j,\;B_{b}=(j+k)\bmod d\bigr{)},$ (27)
then the CGLMP Bell inequality can be stated as
$I_{\operatorname{CGLMP}}\coloneqq$ $\Bigl{[}P(A_{1}=B_{1})+P(B_{1}=A_{2}+1)+P(A_{2}=B_{2})$ (28)
$\quad+P(B_{2}=A_{1})\Bigr{]}$
$\;-\Bigl{[}P(A_{1}=B_{1}-1)+P(B_{1}=A_{2})$
$\quad+P(A_{2}=B_{2}-1)+P(B_{2}=A_{1}-1)\Bigr{]}.$
To extract randomness, we consider the NPO problem from (10) with a single constraint—namely, the violation of the CGLMP inequality by a specific value. It is straightforward to see that the classical bound is $2$, while an outer bound for the quantum value can be readily computed using the NPA hierarchy, which is approximately $2.91485$. In Figure 3 we present our numerical results. Here, the min-entropy is computed from (25) at NPA level $3$, and the bounds from Theorem II.1 are obtained at NPA level $2$ by interchanging the sum and the infimum, using a fine grid (around 30 points). The numerical computations can be completed in seconds to minutes. As shown in Figure 3, the bounds for randomness extracted via Theorem II.1 are strictly higher when using the CGLMP inequality.
222.22.22.42.42.62.62.82.83300.50.5111.51.5CGLMP valueRandomnessmin-entropy (25)Theorem II.1
(a) min-entropy vs. lower bounds from Theorem II.1 for CGLMP for 2233 scenario
4.84.84.854.854.94.94.954.955500.50.511I3322 valueRandomnessmin-entropy (25)Theorem II.1
(b) min-entropy vs. lower bounds from Theorem II.1 for I3322
Figure 3: Lower bounds for the asymptotically extractable randomness $H(A|E)$ in a device independent scenario are shown in dependence of the violation of the Bell inequalities from [35] and the I3322 Bell inequality. The min-entropy values are obtained by optimizing (25) at NPA level $3$, while the direct bounds are obtained with Theorem II.1 and an appropriate fine grid. The behaviour of the CGLMP inequality is qualitatively very similar to the well-known CHSH case, which we have considered in Figure 2. In comparison to that, the I3322 Bell inequality needs a very high violation $\approx 4.83$ (where 44 is the classical, $\approx 5.0035$ the quantum bound) for extraction of randomness. All results can be achieved on personal devices in the amount of minutes.
IV.2 Randomness from I3322
In this section we consider the generalization of the six-state-protocol originally proposed on qubits in [37]. The idea is to consider three PVM’s per party with two outcomes and to generate key out of one of the three bases. In [37] it is shown that the protocol exhibits slightly improved rates if we consider a device dependent scheme on a qubit. To the best of our knowledge, there are no results for fully device independent lower bounds on the conditional von Neumann entropy in dependence of I3322 violation. We provide in Figure 3 bounds for the asymptotically extractable randomness of the I3322 inequality given by
$I_{3322}=$ $\langle A_{1}B_{3}\rangle+\langle A_{2}B_{3}\rangle+\langle A_{3}B_{1}\rangle$ (29)
$+\langle A_{3}B_{2}\rangle+\langle A_{1}B_{2}\rangle+\langle A_{2}B_{1}\rangle-\langle A_{2}B_{2}\rangle$
$-\langle A_{1}B_{1}\rangle+\langle A_{1}\rangle-\langle A_{2}\rangle+\langle B_{1}\rangle-\langle B_{2}\rangle.$
It is straightforward to see that the classical bound is $4$, while an outer bound for the quantum value can be readily computed using the NPA hierarchy, which is approximately $5.0035$. In Figure 3 we present our numerical results. Here, the min-entropy is computed from (25) at NPA level $3$, and the bounds from Theorem II.1 are obtained at NPA level $2$ by interchanging the sum and the infimum, using a fine grid.
V Test on Experimental Data
In this section we report on the combination of our tools with a set of data measured in a real DI-QKD experiment in lab. For this purpose, we performed an analysis of data taken in a early test run of the DI-QKD experiment [7] showcasing that more measurements actually lead to higher certifiably randomness extraction rates in a realistic environment. In order to gain insight into the asymptotically extractable randomness directly from the observed statistics—rather than first estimating a Bell inequality and then optimizing the conditional von Neumann entropy with respect to its violation—several steps are required. Real experimental data inevitably contain small errors and are therefore only approximately non-signalling. However, at the heart of the NPA hierarchy lies the assumption that we work within the commuting operator framework, which inherently imposes a form of non-signalling condition. As a result, simply applying an $\varepsilon$-threshold to the data may yield a bound corresponding to a not-non-signalling distribution, yet this distribution might still be infeasible at a higher level of the NPA hierarchy. For this reason, when aiming to exploit the full informational content of the distribution, we compute the closest non-signalling distribution to the data and enforce it through equality constraints. Since the conditional von Neumann entropy is a continuous function, this approach does introduce a controllable error in the result, while substantially improving numerical stability.
In this section, we present the numerical results of a 2422-experiment, where we have full access to the distribution $p(a,b\mid x,y)$. Table 2 summarizes the outcomes for one-sided randomness extraction. We compare two scenarios for applying Theorem II.1: one uses fewer grid points, and the other interchanges the sum and infimum but employs a large number of grid points. Additionally, we examine two data-extraction strategies: one that uses only the CHSH value from the statistics (approximately $2.5958$), and another that utilizes the complete distribution. Our results show that, up to a precision of $10^{-4}$, there is no significant difference between using only the CHSH value and using the full statistics. In particular we can compare our results to the theoretical result given by [23] which yields for a CHSH violation of $2.5958$ approximately $0.5759$ bits.
A B C
x values
$x=0$ 0.3883 0 0.3915
$x=1$ 0.4178 0 0.4186
$x=2$ 0.6235 0.6235 0.6156
$x=3$ 0.5888 0.5888 0.5779
y values
$y=0$ 0.6072 0.6069 0.5979
$y=1$ 0.6023 0.6023 0.5921
Table 2: In column A we summarize our certified randomness with $22$ grid points in Theorem II.1 but sum and infimum are interchanged and we have used the full statistics as equality constraints. In B we just incorporated the constraints regarding $x=2,3$ and $y=0,1$ which would correspond to statistics to calculate the CHSH value. In C we do not interchange sum and infimum, used the full statistic but use only 77 grid points. All experiments are done at NPA level 22 and include extra monomials of the form $M_{a|x}N_{b|y}P_{k}^{(a)}$. The results are rounded to a precision $\equiv 10^{-4}$.
As a last contribution we report on two-sided randomness extraction whereby we used $22$ grid points in Theorem II.1, in which the sum and infimum operations were interchanged, and full statistical constraints were employed at NPA level $2$ with extra monomials of the form $M_{a|x}N_{b|y}P_{k}^{(a)}$. For the measurement settings $x=1$ and $y=0$, the global randomness value was found to be approximately $0.8520$. Comparing this with the results in Figure 2, yields that it approximately fits exactly what we would expect from CHSH inequality.
VI Comparison to the method of Brown, Fawzi, and Fawzi
In order to demonstrate the potentials our method offers, we will compare it to the method of Brown, Fawzi, and Fawzi (BFF) [22]. Conceptually there are some parallels, both methods build on lower bounding the von Neumann entropy by upper bounding a relative entropy and both methods use integral representation in order to get approximations. The method of Tan, Schwonnek et. al. [21] follows a different ansatz, in this work we will leave a comparison to this method aside. In the device-independent setting, the definition of the conditional von Neumann entropy becomes significantly more challenging, since its standard formulation via matrix logarithms cannot be used. At present, there are essentially two different integral representations of the relative entropy and, in the case of a classical conditioning system, the conditional von Neumann entropy: Kosaki’s formula [38] and Frenkel’s formula [18]. Both formulations provide a way to define the conditional von Neumann entropy in a device-independent setting.
Kosaki’s formula arises from an interpolation result and can be generalized to all operator monotone functions, whereas Frenkel’s formula belongs to the category of $f$-divergences. Since both techniques rely on integration, they can be compared fairly in terms of their efficiency with respect to:
  1. 1.
    the quality of the lower bound,
  2. 2.
    the number of nodes required,
  3. 3.
    runtime in a competitive example, and
  4. 4.
    the size of the moment matrices.
We provide evidence for the comparison with the help of the following examples. The computational experiments were performed on a system equipped with a 13th Gen Intel® Core™ i5-13600KF processor and 128 GB of DDR4 memory operating at a configured speed of 4000 MT/s (maximum rated speed 4800 MT/s).
VI.1 I3322 comparison
4.64.64.84.85500.50.511Bell ScoreEntropy BoundBFF Method, 8 notesThis work, 8 notesThis work, 30 notes
Figure 4: Comparison of entropy bounds obtained using the BFF method and the from this work as a function of the score. On this instance 8 notes is already significanly challening in runtime $\sim 10$s per point. As comparison our method with also 8 notes needs only $\sim 0.5$s (see Figure 5) and gives higher entropies. We also benchmark our method with 30 notes. The runtime is on the order of $\sim 2$s and we can see that the entropy bound further increases.
We investigate randomness extraction from correlations violating the $I_{3322}$ Bell inequality using two approaches: the Brown–Fawzi–Fawzi (BFF) method with relaxations (1), (3), and (4) from [22, Rem. 2.6], and our proposed technique, for which the analogues of [22, Rem. 2.6 (3),(4)] can also be applied. In both methods, we employ $8$ nodes in the quadrature formula for the integration and work at NPA level 2, supplemented with the minimal number of level 3 moment matrix elements required to ensure that the objective functional is well defined. This setup results in moment matrices of size $110\times 110$ for the BFF method and only $62\times 62$ for our technique.
In Figure 4, we present the lower bounds obtained from Theorem II.2 and BFF for both methods as a function of the Bell score. We conclude that they are on a comparable scale. In Figure 5 we provide a runtime comparison: the BFF method requires between approximately $3.58$ s and $18.70$ s per instance over the considered score range, while our method completes in 0.430.430.510.51 s per instance. This demonstrates that, while both methods produce comparable bounds, our approach achieves a substantial speed-up—often exceeding an order of magnitude—due to the significantly smaller moment matrix size resulting from the simpler objective function and the fact that we can assume the operators in Eve’s system to be projective.
4.64.64.84.855055101015152020Bell ScoreRuntime (s)BFF MethodThis work
Figure 5: Runtime comparison of the BFF method and our method (KS) for randomness extraction from the $I_{3322}$ Bell inequality, using 88 quadrature nodes and NPA level 2 with minimal level 3 extensions.
VI.2 CGLMP comparison
222.22.22.42.42.62.62.82.83300.50.5111.51.5Bell ScoreEntropy BoundBFF MethodThis work
Figure 6: Comparison of entropy bounds obtained using the BFF method and the method from this work for the CGLMP inequality as a function of the score.
We investigate randomness extraction from correlations violating the CGLMP Bell inequality using two approaches: the Brown–Fawzi–Fawzi (BFF) method with relaxations (1), (3), and (4) from [22, Rem. 2.6], and our proposed technique, for which the analogues of [22, Rem. 2.6 (3),(4)] can also be applied. In both methods, we employ $8$ nodes in the quadrature formula for the integration and work at NPA level 2, supplemented with the minimal number of level 3 moment matrix elements required to ensure that the objective functional is well defined. This setup results in moment matrices of size $230\times 230$ for the BFF method and only $122\times 122$ for our technique.
In Figure 6, we present the lower bounds obtained from Theorem II.2 and BFF for both methods as a function of the Bell score, demonstrating comparable scales. In Figure 7 we provide a runtime comparison: across the considered score range, the BFF method requires approximately 69.0569.05242.81242.81 s per instance, while our method completes in only 3.083.083.673.67 s per instance. This illustrates a substantial speed-up of our approach, attributable to the significantly smaller moment matrices resulting from a simpler objective functional and the ability to assume projective operators on Eve’s system.
222.22.22.42.42.62.62.82.8330100100200200Bell ScoreRuntime (s)BFF MethodThis work
Figure 7: Runtime comparison of the BFF method and our method (KS) for randomness extraction from the CGLMP inequality, using 88 quadrature nodes and NPA level 2 with minimal level 3 extensions.
VII Summary
This work focuses on optimizing the conditional von-Neumann entropy within a device-independent framework. From a broader perspective, we ask how much randomness can be extracted from two spatially separated, hypothetical “black boxes” located in Alice’s and Bob’s respective labs, where the outcomes follow the statistical laws of quantum theory. Instead of specifying a concrete quantum experiment represented by a Hilbert space $\mathcal{H}$ and a trace-class operator $\rho$ on $\mathcal{H}$, we rely on the achievable statistics defined by general quantum theory. The core principle underlying the presence of genuine randomness here is the monogamy of correlations, which restricts information sharing in non-signalling theories as demonstrated in [11]. This principle remains applicable even with only limited knowledge of the output statistics in (1). Additionally, bounding the capabilities of Alice, Bob, and Eve to the framework of quantum theory is significant not only for applications but also for advancing our understanding of quantum principles. For instance, [39] exploits the gap between general non-signalling theories and quantum theory, highlighting this difference as a powerful tool.
Our approach serves as a versatile framework for deriving lower bounds on the conditional von-Neumann entropy. The main technical contributions of our work include the application of the integral representation from [18] adapted for this setting, as well as the formulation of a non-commutative polynomial optimization problem, which we solve using the NPA hierarchy [19]. This approach bridges the gap between two established methods: those in [22] and [21]. The technique from [22] has become standard in security proofs for device-independent quantum key distribution (DIQKD) and has since been extended to Petz-Rényi divergences [40]. Compared to [22], our method requires fewer NPA variables, as our projections $P_{k}^{(a)}$ in Theorem II.2 reduce complexity relative to the non-Hermitian operators $Z_{a,i}$ in [22]. This reduction streamlines the operator set in the sum-of-squares cone and enables us to impose additional constraints within the kernel of the map $\varphi$ in [30, Fundamental Lemma]. Moreover, our method builds on discretization techniques for integration from [28], while [22] relies on the Gauss-Radau quadrature method from numerical integration theory. However, the advantages of each integration method for the specific computations required remain open for investigation. The method in [21], though distinct from these approaches, recasts the conditional von-Neumann entropy in (2) as
$H(A|Q_{E})_{\rho_{AQ_{E}}}=S(T_{Q_{A}\to A}(\rho_{Q_{A}Q_{B}}))-S(\rho_{Q_{A}Q_{B}}),$ (30)
interpreting it as an entropy production. Here, the channel $T_{Q_{A}\to A}$ acts as the measurement channel on Alice’s system. The method in [21] establishes lower bounds through a variational approach, employing the Gibbs variational principle and Golden-Thompson inequalities. A direct comparison of these conceptually related integral-based methods with the entropy production approach would be a valuable avenue for future research. In section III, we present multiple scenarios to illustrate the effectiveness of our method for randomness extraction. These scenarios include both numerical validations and theoretical comparisons to benchmark results from existing studies.
Firstly, we analyze a $2222$ - scenario (two measurement settings per party, each with two outcomes) under CHSH constraints, which is a standard setting in quantum information theory for testing Bell inequalities. We consider this scenario for both one-sided and two-sided randomness extraction, where randomness is certified by outcomes that are either private to one party (one-sided) or to both parties (two-sided). Our results in this $2222$ - scenario show strong agreement with the analytical values reported in [11], demonstrating the accuracy and robustness of our approach. Additionally, we explore a more intricate $2322$ - scenario, where one party has three measurement settings while the other has two. This scenario is used to examine global randomness extraction, which involves jointly certifying randomness across all outcomes. Our findings, obtained under the statistical framework of Werner states, indicate that this more complex setup does not offer a substantial qualitative advantage over the $2222$ - scenario. This suggests that increasing the measurement settings in this configuration may not necessarily enhance randomness extraction, a result that merits further investigation. An intriguing direction for future research would be to systematically compare different configurations - such as the $2222$, $2322$ and $3322$ - scenarios (where the latter allows each party three measurement settings with two outcomes) - to determine how these variations impact randomness generation under different assumptions of honest implementation or even a random key basis [26]. This could uncover new insights into the optimal design of randomness extraction protocols. Moreover, it would be valuable to examine randomness extraction in more complex non-local games, such as the magic square game discussed in [39]. These games offer alternative frameworks for randomness certification that might provide unique advantages in terms of security and efficiency in quantum protocols.
VIII Acknowledgements
GK thanks Mario Berta, Omar Fawzi, Ian George, Edwin Lobo, Stefano Pironio and Lewis Wooltorton for fruitful discussions. GK thanks Marco Tomamichel for hosting him at CQT Singapore in fall 2024. The author would like to thank Rutvij Bhavsar for insightful discussions and for providing access to the data from [32]. GK acknowledges support from the Excellence Cluster - Matter and Light for Quantum Computing (ML4Q). GK acknowledges funding by the European Research Council (ERC Grant Agreement No. 948139). R.S. is supported by the DFG under Germany’s Excellence Strategy - EXC-2123 QuantumFrontiers - 390837967 and SFB 1227 (DQ-mat) , the Quantum Valley Lower Saxony, and the BMBF projects ATIQ, SEQUIN, QuBRA and CBQD. The authors thank Harald Weinfurter and his research group for measuring the test data that was used for benchmarking our methods.
Appendix A Proof of Theorem II.1
The following is a generalization of [25, Cor. 1] to positive trace-class operators.
Lemma A.1.
Let $\rho,\sigma\in\mathcal{P}(\mathcal{H})$ be positive trace-class operators such that there exist $\mu,\lambda\geq 0$ such that $\mu\,\sigma\leq\rho\leq\lambda\,\sigma$. Then
$D(\rho\|\sigma)=\operatorname{tr}[\rho-\sigma]+\int_{\mu}^{\lambda}\frac{ds}{s}\,\operatorname{tr}^{-}\!\big{[}\rho-s\sigma\big{]}+\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma].$ (31)
Proof.
We start from the integral representation of relative entropy valid for all positive trace-class operators (see [18]):
$D(\rho\|\sigma)=\operatorname{tr}[\rho-\sigma]+\int_{-\infty}^{\infty}\frac{|t|}{(1-t)^{2}}\,\operatorname{tr}^{-}\!\big{[}(1-t)\rho+t\sigma\big{]}\,dt.$ (32)
Split the integral at $t=0$ and $t=1$. For $t\leq 0$ we have 1t>01-t>0 and
$((1-t)\rho+t\sigma)_{-}=(1-t)\bigl{(}\rho-\tfrac{t}{t-1}\sigma\bigr{)}_{-}.$ (33)
With the substitution $s=\tfrac{t}{t-1}$ (which maps $(-\infty,0]\to[0,1]$), one checks that
$\int_{-\infty}^{0}\frac{|t|}{(1-t)^{2}}\operatorname{tr}^{-}[(1-t)\rho+t\sigma]\,dt=\int_{0}^{1}\frac{ds}{s}\,\operatorname{tr}^{-}[\rho-s\sigma].$ (34)
Since $\mu\sigma\leq\rho$, the integrand vanishes for $s\in[0,\mu]$, hence the last integral equals $\int_{\mu}^{1}\frac{ds}{s}\operatorname{tr}^{-}[\rho-s\sigma]$.
For $t\geq 1$, we use
$((1-t)\rho+t\sigma)_{-}=((t-1)\rho-t\sigma)_{+}=(t-1)\bigl{(}\rho-\tfrac{t}{t-1}\sigma\bigr{)}_{+}.$ (35)
With the same substitution $s=\tfrac{t}{t-1}$ (mapping $1,\infty)\to[1,\infty)$),
$\int_{1}^{\infty}\frac{t}{(1-t)^{2}}\operatorname{tr}^{-}[(1-t)\rho+t\sigma]\,dt=\int_{1}^{\infty}\frac{ds}{s}\,\operatorname{tr}^{+}[\rho-s\sigma].$ (36)
Because $\rho\leq\lambda\sigma$, we have $(\rho-s\sigma)_{+}=0$ for all $s\geq\lambda$, so the integral truncates to $\int_{1}^{\lambda}\frac{ds}{s}\operatorname{tr}^{+}[\rho-s\sigma]$.
Insert these two pieces in (32) to get
$D(\rho\|\sigma)=\operatorname{tr}[\rho-\sigma]+\int_{\mu}^{1}\frac{ds}{s}\operatorname{tr}^{-}[\rho-s\sigma]+\int_{1}^{\lambda}\frac{ds}{s}\operatorname{tr}^{+}[\rho-s\sigma].$ (37)
Now use the elementary identity for any self-adjoint $X$: $\operatorname{tr}[X_{+}]=\operatorname{tr}[X_{-}]+\operatorname{tr}[X].$ With $X=\rho-s\sigma$ this gives
$\operatorname{tr}^{+}[\rho-s\sigma]=\operatorname{tr}^{-}[\rho-s\sigma]+\operatorname{tr}[\rho-s\sigma].$ (38)
Hence
$\int_{1}^{\lambda}\frac{ds}{s}\operatorname{tr}^{+}[\rho-s\sigma]=\int_{1}^{\lambda}\frac{ds}{s}\operatorname{tr}^{-}[\rho-s\sigma]+\int_{1}^{\lambda}\frac{ds}{s}\bigl{(}\operatorname{tr}[\rho]-s\,\operatorname{tr}[\sigma]\bigr{)}.$ (39)
Combining with the $\mu,1]$ piece collapses the two $\operatorname{tr}^{-}[\rho-s\sigma]$-integrals into a single integral over $\mu,\lambda]$, and the elementary integral evaluates to
$\int_{1}^{\lambda}\frac{ds}{s}\bigl{(}\operatorname{tr}[\rho]-s\,\operatorname{tr}[\sigma]\bigr{)}=\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma].$ (40)
This yields (31).
Finally, independence of the choice of $\mu,\lambda$ follows from the truncation properties: for $s\leq\mu$ we have $(\rho-s\sigma)_{-}=0$, and for $s\geq\lambda$ we have $(\rho-s\sigma)_{+}=0$. Equivalently, if $\lambda^{\prime}\geq\lambda$ then on $(\lambda,\lambda^{\prime}]$ one has $(\rho-s\sigma)_{-}=s\,\operatorname{tr}[\sigma]-\operatorname{tr}[\rho]$, so the change in $\int\frac{ds}{s}\operatorname{tr}^{-}[\rho-s\sigma]$ is exactly canceled by the change in $\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma]$, and an analogous statement holds at the lower limit $\mu$. ∎
Proof of Theorem II.1:
Proof.
Fix a partition $0<t_{1}<\cdots<t_{r}$ with $t_{1}=\mu$ and $t_{r}=\lambda$. Following the tools in [28], set
$y_{k}:=\sup_{0\leq P\leq\mathds{1}}\operatorname{tr}\big{[}P(\sigma t_{k}-\rho)\big{]}.$ (41)
Then [28] yields the estimate
$\int_{\mu}^{\lambda}\frac{ds}{s}\,\operatorname{tr}^{+}[\sigma s-\rho]$ $\leq y_{1}\Big{[}\Big{(}1+\frac{t_{1}}{t_{2}-t_{1}}\Big{)}\ln\frac{t_{2}}{t_{1}}-1\Big{]}+y_{r}\Big{[}1-\frac{t_{r-1}}{t_{r}-t_{r-1}}\ln\frac{t_{r}}{t_{r-1}}\Big{]}$ (42)
$\qquad+\sum_{k=2}^{r-1}y_{k}\Big{[}\Big{(}1+\frac{t_{k}}{t_{k+1}-t_{k}}\Big{)}\ln\frac{t_{k+1}}{t_{k}}-\frac{t_{k-1}}{t_{k}-t_{k-1}}\ln\frac{t_{k}}{t_{k-1}}\Big{]}.$
It is convenient to encode the coefficients by
$\alpha_{k}:=\begin{cases}-\Big{[}\Big{(}1+\dfrac{t_{1}}{t_{2}-t_{1}}\Big{)}\ln\dfrac{t_{2}}{t_{1}}-1\Big{]}&k=1,\\[4.0pt] -\Big{[}1-\dfrac{t_{r-1}}{t_{r}-t_{r-1}}\ln\dfrac{t_{r}}{t_{r-1}}\Big{]}&k=r,\\[6.0pt] -\Big{[}\Big{(}1+\dfrac{t_{k}}{t_{k+1}-t_{k}}\Big{)}\ln\dfrac{t_{k+1}}{t_{k}}-\dfrac{t_{k-1}}{t_{k}-t_{k-1}}\ln\dfrac{t_{k}}{t_{k-1}}\Big{]}&\text{else},\end{cases}$
and
$\beta_{k}:=\begin{cases}\Big{[}\Big{(}1+\dfrac{t_{1}}{t_{2}-t_{1}}\Big{)}\ln\dfrac{t_{2}}{t_{1}}-1\Big{]}\,t_{1}&k=1,\\[4.0pt] \Big{[}1-\dfrac{t_{r-1}}{t_{r}-t_{r-1}}\ln\dfrac{t_{r}}{t_{r-1}}\Big{]}\,t_{r}&k=r,\\[6.0pt] \Big{[}\Big{(}1+\dfrac{t_{k}}{t_{k+1}-t_{k}}\Big{)}\ln\dfrac{t_{k+1}}{t_{k}}-\dfrac{t_{k-1}}{t_{k}-t_{k-1}}\ln\dfrac{t_{k}}{t_{k-1}}\Big{]}\,t_{k}&\text{else}.\end{cases}$
For the interval $0,\mu]$, use the convexity of $s\mapsto\operatorname{tr}^{+}[\sigma s-\rho]$ [28, Lem. 1] and the fact that $\operatorname{tr}^{+}[\sigma s-\rho]\to 0$ as $s\to 0$ for positive $\sigma,\rho\in\mathcal{P}(\mathcal{H})$. For $s\in[0,\mu]$,
$\operatorname{tr}^{+}[\sigma s-\rho]\leq\Big{(}1-\frac{s}{\mu}\Big{)}\operatorname{tr}^{+}[\sigma\cdot 0-\rho]+\frac{s}{\mu}\operatorname{tr}^{+}[\sigma\mu-\rho]=\frac{s}{\mu}\operatorname{tr}^{+}[\sigma\mu-\rho].$
Hence
$\int_{0}^{\mu}\frac{ds}{s}\,\operatorname{tr}^{+}[\sigma s-\rho]$ $\leq\operatorname{tr}^{+}[\sigma\mu-\rho]=\sup_{0\leq P\leq\mathds{1}}\operatorname{tr}\big{[}P(\sigma\mu-\rho)\big{]},$
which corresponds to the choice $\alpha_{0}=-1$ and $\beta_{0}=\mu$.
Inserting (41) into (42), and using the definitions of $\alpha_{k},\beta_{k}$ for $0\leq k\leq r$, we obtain with A.1
$D(\rho\|\sigma)\leq\frac{1}{\ln 2}\Big{(}\operatorname{tr}[\rho-\sigma]+\sup_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\sum_{k=0}^{r}\operatorname{tr}\big{[}P_{k}(\alpha_{k}\rho+\beta_{k}\sigma)\big{]}+\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma]\Big{)},$
which is exactly (13). This proves the claim. ∎
Appendix B Proof Theorem II.2
Proof.
For simplicity we neglect all the way the whole constraints regadarding Bell-type expressions in (16). Those constraints are well-known and came along many times in similar calculations regarding the device independent bounds on Bell violations. We start with the following expression, which is basically the step in (15), which is certified by Theorem II.1 and $\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}})$
$\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}H(A|X=\tilde{x},Q_{E})_{\rho_{AQ_{E}|\tilde{x}}}$= $\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}-D(\rho_{AQ_{E}|\tilde{x}}\|\mathds{1}_{A}\otimes\rho_{Q_{E}})$ (43)
$\geq$ $\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\frac{1}{\ln 2}(\operatorname{tr}[\mathds{1}_{A}\otimes\rho_{Q_{E}}-\rho_{AQ_{E}|\tilde{x}}$]
$\quad+\sum_{k=0}^{r}\operatorname{tr}[P_{k}(-\alpha_{k}\rho_{AQ_{E}|\tilde{x}}-\beta_{k}\mathds{1}_{A}\otimes\rho_{Q_{E}})].$
In a next step, we use the fact that
$\operatorname{tr}[\mathds{1}_{A}\otimes\rho_{Q_{E}}-\rho_{AQ_{E}|\tilde{x}}]=|A|-1$ (44)
and that each of the $P_{k}$ is a projector in a classical quantum system such that we can write
$P_{k}=\sum_{a\in A}|a\rangle\langle a|\otimes P_{k}^{(a)}.$ (45)
For some positive operator $P_{k}^{(a)}$. Calculating $(P_{k})^{2}$ then yields that even the $P_{k}^{(a)}$ have to be projections if $P_{k}$ is a projection. In addition we use the following calculation for $\rho_{AQ_{E}|\tilde{x}}$
$\rho_{AQ_{E}|\tilde{x}}$ $=\sum_{a\in A}p(a|x)|a\rangle\langle a|\otimes\rho_{Q_{E}|A=a,\tilde{x}}$ (46)
$=\sum_{a\in A}|a\rangle\langle a|\otimes\operatorname{tr}_{Q_{A}Q_{B}}(\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes\mathds{1}_{B}).$
Before we proceed we briefly argue that of course Eve’s marginal does not depend on $\tilde{x}$. This is the non-signalling property of quantum theory or one may calculate it directly
$\rho_{Q_{E}|\tilde{x}}$ $=\operatorname{tr}_{A}[\rho_{AQ_{E}|\tilde{x}}$] (47)
$=\operatorname{tr}_{A}[\sum_{a\in A}p(a|x)|a\rangle\langle a|\otimes\rho_{Q_{E}|A=a,\tilde{x}}$]
$=\sum_{a\in A}\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes\mathds{1}_{Q_{B}}]$
$=\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}\sum_{a\in A}M_{a|\tilde{x}}\otimes\mathds{1}_{Q_{B}}]$
$=\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}\mathds{1}_{Q_{A}}\otimes\mathds{1}_{Q_{B}}]$
$=\rho_{Q_{E}}.$
Now we insert (44),(46) and (47) into (43) to get
$\ldots$ $=\frac{1}{\ln 2}\big{(}|A|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{a},\ldots,P_{r}^{a}}\sum_{k=0}^{r}-\alpha_{k}\operatorname{tr}[\sum_{a^{\prime}\in A}|a^{\prime}\rangle\langle a^{\prime}|\otimes\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes\mathds{1}_{Q_{B}Q_{E}}]\sum_{a\in A}|a\rangle\langle a|\otimes P_{k}^{(a)}]$ (48)
$\hskip 170.71652pt+\sum_{k=0}^{r}-\beta_{k}\operatorname{tr}[\mathds{1}_{A}\otimes\rho_{Q_{E}}\sum_{a\in A}|a\rangle\langle a|\otimes P_{k}^{(a)}]\big{)}$
$=\frac{1}{\ln 2}\big{(}|A|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{a},\ldots,P_{r}^{a}}\sum_{k=0}^{r}\sum_{a\in A}-\alpha_{k}\operatorname{tr}[|a\rangle\langle a|\otimes\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes\mathds{1}_{Q_{B}Q_{E}}]P_{k}^{(a)}]$
$\hskip 170.71652pt+\sum_{k=0}^{r}\sum_{a\in A}-\beta_{k}\operatorname{tr}[|a\rangle\langle a|\otimes\rho_{Q_{E}}P_{k}^{(a)}]\big{)}$
$=\frac{1}{\ln 2}\big{(}|A|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{a},\ldots,P_{r}^{a}}\sum_{k=0}^{r}\sum_{a\in A}-\alpha_{k}\operatorname{tr}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes\mathds{1}_{Q_{B}}\otimes P_{k}^{(a)}]-\beta_{k}\operatorname{tr}[\rho_{Q_{A}Q_{B}Q_{E}}\mathds{1}_{Q_{A}}\otimes\mathds{1}_{Q_{B}}\otimes P_{k}^{(a)}]\big{)}.$
Now using the fact that $\rho_{AQ_{E}|\tilde{x}}\leq\mathds{1}_{A}\otimes\rho_{Q_{E}}$ (due to the fact that the state is classical quantum) and thus $\lambda=1$ yields $\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma]=0$. Moreover, as discussed in Theorem II.1 and the proof in Appendix A, $\mu=0$ can be handled without problems such that the constraints coming from the integral representation in (6) become trivial. Moreover, the operators $P_{k}^{(a)}\in\mathcal{B}(\mathcal{H}_{Q_{E}})$ such that they commute with all local measurements on Alice and Bob’s side respectively. Adding the fact that they are projections from (45), we conclude all the constraints stated in Theorem II.2. Observing the purity argument of $\rho_{Q_{A}Q_{B}Q_{E}}$ by Naimark’s Theorem yields the assertion of the Theorem II.2, if we replace tensor products with commuting operators. ∎
Appendix C Global randomness from the full distribution
We show similarly to Appendix B how to get lower bounds for $H(AB|X=\tilde{x},Y=\tilde{y},Q_{E})$ and abbreviate steps similar to Appendix B. Similarly, we neglect all Bell-type expressions in (16). Applying Theorem II.1 for $\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}})$ yields
$\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}H(AB|X=\tilde{x},Y=\tilde{y}Q_{E})_{\rho_{AQ_{E}|\tilde{x}}}$= $\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}-D(\rho_{ABQ_{E}|\tilde{x},\tilde{y}}\|\mathds{1}_{AB}\otimes\rho_{Q_{E}})$ (49)
$\geq$ $\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\frac{1}{\ln 2}(\operatorname{tr}[\mathds{1}_{AB}\otimes\rho_{Q_{E}}-\rho_{ABQ_{E}|\tilde{x},\tilde{y}}$]
$\quad+\sum_{k=0}^{r}\operatorname{tr}[P_{k}(-\alpha_{k}\rho_{ABQ_{E}|\tilde{x},\tilde{y}}-\beta_{k}\mathds{1}_{AB}\otimes\rho_{Q_{E}})].$
Using the fact that
$\operatorname{tr}[\mathds{1}_{AB}\otimes\rho_{Q_{E}}-\rho_{ABQ_{E}|\tilde{x},\tilde{y}}]=|A|\cdot|B|-1$ (50)
and that each of the $P_{k}$ is a projector in a classical quantum system, we rewrite
$P_{k}=\sum_{a\in A,b\in B}|a,b\rangle\langle a,b|\otimes P_{k}^{(ab)}.$ (51)
with positive operators $P_{k}^{(ab)}\in\mathcal{P}(\mathcal{H}_{E})$. Furthermore, we can apply the following calculations for $\rho_{ABQ_{E}|\tilde{x},\tilde{y}}$
$\rho_{ABQ_{E}|\tilde{x},\tilde{y}}$ $=\sum_{a\in A,b\in B}p(a,b|x,y)|a,b\rangle\langle a,b|\otimes\rho_{Q_{E}|A=a,B=b,\tilde{x},\tilde{y}}$ (52)
[…]
Now we insert (50) and (52) into (49) to get
$\ldots$ $=\frac{1}{\ln 2}\big{(}|A|\cdot|B|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{a},\ldots,P_{r}^{a}}\sum_{k=0}^{r}-\alpha_{k}\operatorname{tr}[\sum_{a^{\prime}\in A,b^{\prime}\in B}|a^{\prime},b^{\prime}\rangle\langle a^{\prime},b^{\prime}|\otimes\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes N_{b|\tilde{y}}\otimes\mathds{1}_{Q_{E}}]$ (53)
$\hskip 85.35826pt\sum_{a\in A,b\in B}|a,b\rangle\langle a,b|\otimes P_{k}^{(ab)}]+\sum_{k=0}^{r}-\beta_{k}\operatorname{tr}[\mathds{1}_{AB}\otimes\rho_{Q_{E}}\sum_{a\in A,b\in B}|a,b\rangle\langle a,b|\otimes P_{k}^{(ab)}]\big{)}$
$=\frac{1}{\ln 2}\big{(}|A|\cdot|B|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{ab},\ldots,P_{r}^{ab}}\sum_{k=0}^{r}\sum_{a\in A,b\in B}-\alpha_{k}\operatorname{tr}[|a,b\rangle\langle a,b|\otimes\operatorname{tr}_{Q_{A}Q_{B}}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes N_{b|\tilde{y}}\otimes\mathds{1}_{Q_{E}}]P_{k}^{(ab)}]$
$\hskip 170.71652pt+\sum_{k=0}^{r}\sum_{a\in A,b\in B}-\beta_{k}\operatorname{tr}[|a,b\rangle\langle a,b|\otimes\rho_{Q_{E}}P_{k}^{(ab)}]\big{)}$
$=\frac{1}{\ln 2}\big{(}|A|\cdot|B|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{ab},\ldots,P_{r}^{ab}}\sum_{k=0}^{r}\sum_{a\in A,b\in B}-\alpha_{k}\operatorname{tr}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes N_{b|\tilde{y}}\otimes P_{k}^{(ab)}]$
$\hskip 113.81102pt-\beta_{k}\operatorname{tr}[\rho_{Q_{A}Q_{B}Q_{E}}\mathds{1}_{Q_{A}Q_{B}}\otimes P_{k}^{(ab)}]\big{)}.$
Now using the fact that $\rho_{ABQ_{E}|\tilde{x},\tilde{y}}\leq\mathds{1}_{AB}\otimes\rho_{Q_{E}}$ (due to the fact that the state is classical quantum) and thus $\lambda=1$ yields $\ln\lambda+1-\lambda=0$. Adding all arguments from Appendix B we deduce from Appendix B a lower bound on two-sided global randomness
$H(AB|X=0,Y=0,Q_{E})$ $\geq\frac{1}{\ln 2}\big{(}|A|+|B|-1+\inf_{\rho_{Q_{A}Q_{B}Q_{E}}}\inf_{P_{0}^{a},\ldots,P_{r}^{a}}\sum_{k=0}^{r}\sum_{a\in A,b\in B}-\alpha_{k}\operatorname{tr}[\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes N_{b|\tilde{y}}\otimes P_{k}^{(ab)}]$ (54)
$\hskip 85.35826pt-\beta_{k}\operatorname{tr}[\rho_{Q_{A}Q_{B}Q_{E}}\mathds{1}_{Q_{A}Q_{B}}\otimes P_{k}^{(ab)}]\big{)}.$ (55)
Appendix D Algebraic Perspective on the NPA hierarchy
The formulation presented in Theorem II.2 uses the language of Hilbert spaces and optimization, establishing a framework based on all possible Hilbert space representations for the physical system under consideration. However, for solving the optimization problem more effectively, it is advantageous to recast the problem within the framework of a $C^{\star}$-algebra. To this end, we define the complex free $\star$-algebra, denoted $\mathcal{F}(\mathcal{G})$, generated by a set of symbols $\mathcal{G}$ as described in [20]. In our case, the generating set $\mathcal{G}$ is given by
$\mathcal{G}\coloneqq\{P_{k}^{a},M_{a|x},N_{b|y}\ |\ 0\leq k\leq r,\ x\in X,\ y\in Y,\ a\in A,\ b\in B\}.$ (56)
According to [20], we can impose a norm on the algebra $\mathcal{F}(\mathcal{G}|\mathcal{R})$ for a set of relations $\mathcal{R}$. This process allows us, via a completion procedure, to obtain a $C^{\star}$-algebra denoted $C^{\star}(\mathcal{G}|\mathcal{R})$. Lemma 4 in [20] establishes that there is an equivalence between this algebraic formulation and an optimization over its representations.
Within a $C^{\star}$-algebra framework, inner approximations of the positive cone are naturally represented as follows:
$\Sigma_{2}^{(1)}\subset\ldots\subset\Sigma_{2}^{(n)}\subset\Sigma_{2}^{(n+1)}\subset\ldots\subset C^{\star}(\mathcal{G}|\mathcal{R})^{+}.$ (57)
Here, the sequence $\{\Sigma_{2}^{(n)}\}$ represents nested inner approximations of the positive cone $C^{\star}(\mathcal{G}|\mathcal{R})^{+}$. Moving to the dual cones of each $\Sigma_{2}^{(n)}$, we arrive at a corresponding sequence:
$(\Sigma_{2}^{(1)})^{\star}\supset\ldots\supset(\Sigma_{2}^{(n)})^{\star}\supset(\Sigma_{2}^{(n+1)})^{\star}\supset\ldots\supset\mathcal{P}(C^{\star}(\mathcal{G}|\mathcal{R})),$ (58)
where $C^{\star}(\mathcal{G}|\mathcal{R})^{+}$ denotes the natural positive cone of the $C^{\star}$-algebra $C^{\star}(\mathcal{G}|\mathcal{R})$, and $\mathcal{P}(C^{\star}(\mathcal{G}|\mathcal{R}))$ denotes the space of positive, linear, and continuous functionals on $C^{\star}(\mathcal{G}|\mathcal{R})$. By choosing appropriate subsets $\Sigma_{2}^{(n)}$, we can ensure that the dual cone $(\Sigma_{2}^{(n)})^{\star}$ is representable as a semidefinite program (SDP), facilitating a two-step solution approach.
First, we formulate the initial optimization problem as follows:
$c^{\star}\coloneqq\inf$ $\omega(F_{0})$ (59)
s.t. $\omega(F_{i})\leq f_{i}\quad 1\leq i\leq m$
$\omega\in\mathcal{P}(C^{\star}(\mathcal{G}|\mathcal{R})).$
In the first relaxation, we approximate the problem by replacing $\mathcal{P}(C^{\star}(\mathcal{G}|\mathcal{R}))$ with the dual cone $(\Sigma_{2}^{(n)})^{\star}$, yielding
$c^{\star}\geq\inf$ $\omega(F_{0})$ (60)
s.t. $\omega(F_{i})\leq f_{i}\quad 1\leq i\leq m$
$\omega\in(\Sigma_{2}^{(n)})^{\star}.$
In the second step, we exploit the SDP-representability of $(\Sigma_{2}^{(n)})^{\star}$ within a finite-dimensional matrix algebra, constructed via a set of operators $\{K_{j}\}$. By mapping each condition $\{F_{i}\}$ to specific matrices in this SDP framework, we arrive at the final SDP formulation:
$\inf$ $\operatorname{tr}(\rho M_{F_{0}})$ (61)
s.t. $\operatorname{tr}(\rho M_{F_{i}})\geq f_{i}\quad 1\leq i\leq m$
$\operatorname{tr}(\rho K_{j})=0\quad 1\leq j\leq n$
$\rho\geq 0.$
An illustrative example of this methodology is the sum-of-squares (SOS) cones, which were initially introduced in [19]. For SOS cones, the Python package described in [31] automates the translation of generators and relations into an SDP within a matrix algebra framework.
Appendix E Additional discussion for the numerics
We repeat in this section the resulting optimization problem from Theorem II.2
$H(A|X=\tilde{x},Q_{E})\geq\inf$ $\frac{1}{\ln 2}(|A|-1+\sum_{k=0}^{r}\sum_{a\in A}\langle\psi_{Q_{A}Q_{B}Q_{E}}|-\alpha_{k}M_{a|x}P_{k}^{(a)}-\beta_{k}P_{k}^{(a)}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle)$ (62)
$\sum_{abxyi}c_{abxyi}\langle\psi_{Q_{A}Q_{B}Q_{E}}|M_{a|x}N_{b|y}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle\geq q_{i}\quad 1\leq i\leq m$
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X}$
$\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X}$
$N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
$[M_{a|x},N_{b|y}]=[M_{a|x},P_{k}^{(a)}]=[N_{b|y},P_{k}^{(a)}]=0\quad b\in B,\ a\in A,\ x\in\mathcal{X},\ y\in\mathcal{Y},0\leq k\leq r$
$(P_{k}^{(a)})^{2}=P_{k}^{(a)},\quad(P_{k}^{(a)})^{\star}=P_{k}^{(a)}\quad 1\leq k\leq r,\ a\in A.$
As discussed in Appendix D, the relations
$\sum_{a}M_{a|x}=\mathds{1}_{A},\quad x\in\mathcal{X}$ (63)
$\sum_{b}N_{b|y}=\mathds{1}_{B},\quad y\in\mathcal{Y}$
$M_{a|x}\geq 0\quad a\in A,\ x\in\mathcal{X}$
$N_{b|y}\geq 0\quad b\in B,\ y\in\mathcal{Y}$
$[M_{a|x},N_{b|y}]=[M_{a|x},P_{k}^{(a)}]=[N_{b|y},P_{k}^{(a)}]=0\quad b\in B,\ a\in A,\ x\in\mathcal{X},\ y\in\mathcal{Y},0\leq k\leq r$
$(P_{k}^{(a)})^{2}=P_{k}^{(a)},\quad(P_{k}^{(a)})^{\star}=P_{k}^{(a)}\quad 1\leq k\leq r,\ a\in A.$
become equality conditions in (61) expressed within the $\{K_{j}\}$. We are now left with the objective function
$\frac{1}{\ln 2}\left(|A|-1+\sum_{k=0}^{r}\sum_{a\in A}\langle\psi_{Q_{A}Q_{B}Q_{E}}|-\alpha_{k}M_{a|x}P_{k}^{(a)}-\beta_{k}P_{k}^{(a)}|\psi_{Q_{A}Q_{B}Q_{E}}\rangle\right),$ (64)
along with some associated constraints. As discussed in [22, Rem. 2.6], we can implement several optimizations to accelerate the numerical evaluations:
  1. 1.
    Modifying the infimum and one of the summations in (64) provides lower bounds, reducing the number of required operators $P_{k}^{(a)}$ to just $\|A\|$. This adjustment significantly accelerates the numerical computations, allowing for the use of a grid with any desired precision.
  2. 2.
    We expand the generating set for the sum-of-squares cone and NPA level $2$ by incorporating additional monomials. In alignment with the approach in [22], we include elements of the form $P_{k}M_{a|x}N_{b|y}$, which enriches the algebraic structure and improves convergence properties.
  3. 3.
    Finally, we can set the NPA matrices to be real-valued, which simplifies computations further.
Applying these optimizations enables all computations to complete within seconds.